### v1

# Task Facet Learning: A Structured Approach to Prompt Optimization

Abstract
Given a task in the form of a basic description and its training examples, prompt optimization is the problem of synthesizing the given information into a text prompt for a large language model (LLM). Humans solve this problem by also considering the different facets that define a task (e.g., counter-examples, explanations, analogies) and including them in the prompt. However, it is unclear whether existing algorithmic approaches, based on iteratively editing a given prompt or automatically selecting a few in-context examples, can cover the multiple facets required to solve a complex task. In this work, we view prompt optimization as that of learning multiple facets of a task from a set of training examples. We identify and exploit structure in the prompt optimization problem — first, we find that prompts can be broken down into loosely coupled semantic sections that have a relatively independent effect on the prompt’s performance; second, we cluster the input space and use clustered batches so that the optimization procedure can learn the different facets of a task across batches. The resulting algorithm, UniPrompt, consists of a generative model to generate initial candidates for each prompt section; and a feedback mechanism that aggregates suggested edits from multiple mini-batches into a conceptual description for the section. Empirical evaluation on multiple datasets and a real-world task shows that prompts generated using UniPrompt obtain higher accuracy than human-tuned prompts and those from state-of-the-art methods. In particular, our algorithm can generate long, complex prompts that existing methods are unable to generate. Code for UniPrompt will be available at https://aka.ms/uniprompt.
1 Introduction
Given a task, choosing an input prompt is a key part of optimizing Large Language Model’s (LLM) performance [14, 31]. Minor changes in prompt can lead to performance gains or losses, necessitating prompt engineering [16] for effective model utilization. Typically, manually-developed prompts combine task description with a few in-context examples, along with modifiers like chain-of-thought [14]. For greater accuracy, human prompt engineers spend considerable time to identify errors with a current prompt, consider the different facets of a task (e.g., counter-examples, explanations, analogies) that may fix those errors, and include them in the prompt. For instance, for a hate speech classification task, in addition to the definition, it may be helpful to specify the facets that lead to hate speech: the context of conversation, identifying intent, and differentiating hate speech from opinions or closely-related concepts such as vulgarity and profanity.
To avoid the cumbersome process of manually creating prompts, recent work aims to automate the process of generating natural language prompts that are also interpretable. Since language tokens are discrete, this leads to a challenging discrete optimization problem with a combinatorial space of possible outputs. Techniques for prompt optimization can be divided in two categories: non-directional, e.g., random search [36, 35] and genetic algorithms [31, 7], where the sampling of new input is “random” and does not explicitly aim to reduce error on a train set; and directional, where the sampling of new input depends on some error measure on a representative train sample. Recently, more complex methods are proposed in the second category including reinforcement learning [33, 6] and updating prompts using auxiliary LLMs to reduce incorrect predictions over a train sample [12, 22]. While all these techniques focus on editing, adding, or deleting parts of a given prompt, they are developed with the goal of obtaining a concise description of the task. None of them focus on ensuring that multiple facets of a task are added to the prompt.
In this paper, we study how to optimize prompts to cover diverse, multiple facets of a task and improve overall accuracy. To develop a suitable algorithm, we analyze properties of the prompt optimization problem. First, we check how sensitive LLMs are to small changes in the input prompt, since directional optimization algorithms may not be as useful if LLMs are highly sensitive. Using a notion of probabilistic Lipschitz continuity, we find that LLMs become less sensitive with size. Across prompt edits with the same semantic meaning, Llama-13B accuracy fluctuates up to 63% of the minimum accuracy whereas the relative accuracy change in GPT-3.5 and GPT-4 models is 6% and 4% respectively. As a result, probabilistic Lipschitz constant for GPT-3.5 and GPT-4 is significantly smaller than that for Llama-13B, indicating that directional optimization methods may be best suited for larger models. Second, we study how LLM accuracy changes as more information is added to the prompt. Ideally, we would want a diminishing marginal returns (sub-modularity) property so that greedy addition of different facets can lead to a reasonable approximation of the true optimum. We find that diminishing returns property largely holds when adding different facets as new sections to the prompt, but it does not hold when directly adding in-context examples.
Our analysis provides two insights: 1) directional optimization may be feasible for models like GPT-3.5 having a low probabilistic Lipschitz constant; 2) including facets as additional text sections may lead to a well-behaved optimization problem. Based on these, we propose an algorithm, UniPrompt, that can generate the full prompt from scratch, given a one line description for a task. Specifically, we identify a prompt structure that provides different task facets as sections in the prompt and initializes them with text generated from a language model. To develop a high-accuracy prompt, we use an auxiliary LLM to obtain feedback about example predictions with the current prompt and decide to make edits at a section level (add, edit or delete) in each iteration. Our key insight is that feedback on a single example or a randomly selected batch of examples does not yield generalizable facet descriptions. Instead, we propose generating edits at a minibatch level and aggregating the edits at the batch level to yield a single edit that conveys a broad concept relevant to the task. To further increase the chances of facet learning, we cluster the input examples beforehand and use clustered mini-batches (see Figure 1).
We evaluate UniPrompt on zero-shot (no in-context examples) and few-shot settings on four different tasks. In the zero-shot setting, UniPrompt consistently achieves higher accuracy than existing prompt optimization methods. On Ethos, a hate speech dataset, UniPrompt obtains 94% accuracy whereas the next best method obtains 82%. In the few-shot setting, we compare UniPrompt to MedPrompt [21], a state-of-the-art prompt composition method. We find that UniPrompt, requiring only one LLM call at inference time, obtains the same accuracy as MedPrompt that requires five calls. If we allow multiple calls to UniPrompt, we obtain over 4% accuracy gains. We also evaluate UniPrompt on a real-world semantic matching task in a web search engine. Compared to the best manual prompt, the prompt generated from UniPrompt leads to over 5% increase in accuracy on the negative class and nearly 2% accuracy increase overall.
Figure 1: Existing prompt optimization methods (left) versus UniPrompt (right) on the SciQ [29] dataset: [Left] State-of-the-art prompt optimization methods such as ProTeGi  [22] sample from the questions wrongly answered by the current prompt, and use an expert LLM (such as GPT-4) to obtain feedback on the mistakes made; This approach tends to overfit to specific examples, as can been seen from the proposed prompts $p^{\prime}$; leading to limited generalization. [Right] In contrast, UniPrompt identifies key task facets that generalize across examples by: (1) clustering examples into groups with similar task facets, and (2) employing a two-tier feedback-based update strategy. The resulting prompt updates extract generalizable concepts from the specific examples seen.
2 Characterizing the Prompt Optimization Problem
Assume access to black-box solver LLM $f:\mathcal{X}\rightarrow\mathbb{R}$ that takes as input a prompt $\mathbf{x}\in\mathcal{X}$ and outputs the average accuracy on a validation set $D_{v}$. Since the set of prompts is combinatorially large, we assume that all prompts can be embedded in a fixed $p$-dimensional space such that distance between two prompts in the vector space correspond to their semantic similarity. Thus, each prompt $\mathbf{x}$ can be expressed as a vector in $\mathbb{R}^{p}$. The prompt optimization problem can be written as $\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x};D_{v})$.
[…]
2.1 Is directional optimization feasible?
Previous work has shown that LLMs can be brittle to their input: changing the prompt slightly can create a significant difference in performance [37]. To understand whether optimization algorithms may work, it is important to understand how the surface of $f$ looks. If $f$ is brittle everywhere to small changes, then random search over prompts (such as in [36]) may be a competitive method; it is unclear whether optimization can help. On the other hand, if the change in $f$ given small changes to the input prompt is bounded, then optimization algorithms can hope to perform significantly better than random search.
More formally, we ask whether the prompt optimization problem is well-conditioned. Typically, conditioning can be determined by the Hessian. However, since $f$ is assumed to be black-box, we approximate it by measuring sensitivity, or more specifically, Lipschitz continuity near the optimal solution. Based on prior work on defining continuity of neural networks [18], we use a probabilistic notion.
Definition 1
Probabilistic Lipschitz Continuity [18]. Given a probability distribution over inputs $\mathcal{X}$, $r\geq 0$, a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(L,\epsilon)$-probabilistically Lipschitz with constant $L\geq 0$, if
$\Pr_{\mathbf{x},\mathbf{x}^{\prime}\sim\mathcal{X}}[\operatorname{d}(f(\mathbf{x}),f(\mathbf{x}^{\prime}))\leq L.\operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\ |\ \operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\leq r]\geq 1-\epsilon$ (1)
where $d$ is a distance measure such as $\ell_{1}$ or $\ell_{2}$ norm.
Note the focus on small changes in input through the parameter $r$. Intuitively, the Lipschitz property bounds the maximum change in $f$ given a small change in input prompt. Typically, the lower bound of error for any sequential optimization algorithm over $f$ is directly proportional to the Lipschitz constant $L$ [17]. Therefore, for faster convergence, it is desirable to have a low $L$, especially near the optimal solution. Empirically, we estimate $L$ by sampling task-relevant prompts so that they are close to the optimal solution. Then we make small changes to the prompt such that the semantic meaning stays the same and measure the change in $f$. We show the change in $f$ per change in input for GPT-4, GPT-3.5, and Llama2-13B models in Figure 2(a) for the Ethos dataset commonly used in prompt optimization [22, 19, 32, 23, 10]. Assuming $\epsilon=0.05$, probabilistic Lipschitz constant $L$ for GPT-3.5 and GPT-4 is $<1$, whereas it is over three times more for Llama2-13B. Thus, as the size of the model increases, the probabilistic Lipschitz constant decreases, indicating that larger models like GPT-3.5 and GPT-4 are more robust to small changes and thus amenable to optimization algorithms. For Llama, the absolute change in accuracy ranges between 0.33 to 0.54, indicating a 63% change in accuracy even across prompts with the same semantic meaning. In contrast, the relative change in accuracy for GPT-3.5 and GPT-4 is 6% and 4% respectively. (See Appendix A.1 for experimental details)
Figure 2: On the Ethos dataset, (a) estimating (probabilistic) Lipschitz constant of models (Definition 1); (b) diminishing marginal utility (submodularity) of GPT-3.5 on two types of prompting strategies.
2.2 Can higher-order abstractions of task examples help prompt optimization?
Consider in-context learning [3], where the prompt $\mathbf{x}$ is composed of a simple task description and a set of labeled examples from the task. The optimal prompt should consider all the aspects of the task which can only be discovered by exhaustively looking at the examples. But, in practice, we are constrained by (a) availability of a limited number of training examples, (b) prompt size. It is beneficial for in-context learning to consider the marginal utility of examples [34, 8]. For instance, it would help to add examples where the model fails rather than ones where the model already succeeds (see Table 4 in Appendix). This suggests that one could use a greedy algorithm for iteratively optimizing the prompt by finding failing examples and adding to the prompt. But would such a procedure converge to a good solution?
We can view $f(\mathbf{x})$ as a set function over the examples (that compose $\mathbf{x}$). A natural property to examine is submodularity (where greedy optimization provably converges to a good solution under some conditions [15]), i.e., the property of diminishing marginal utility of examples. The marginal utility of adding an example $e$ to the few-shot prompt corresponding to a set of examples $A$ is $\Delta_{A}=f(A\cup\{e\})-f(A)$. Let $B\subset A$ denote a subset of $A$. We define $f_{\text{fewshot}}$ to be the validation accuracy of GPT-3.5 using few-shot prompts on the Ethos dataset (see Section 4). In Figure 2 (b), we show the distribution of diminishing marginal utilities for $f_{\text{fewshot}}$ (light blue), i.e., we sample multiple $A,B$ sets of examples from the dataset and compute $\Delta_{B}-\Delta_{A}$, which is expected to be non-negative for submodular functions. As we see, the marginal utility differences are nearly uniformly random in $[-0.01,0.01]$, indicating that there is no submodularity structure to $f_{\text{fewshot}}$.
A recurring theme [1, 2] in the prompts that humans engineer is to organize task information in “sections” (introduction, context, etc.), treating prompt like a structured document. When humans engineer a prompt, they do it systematically by editing specific sections of the prompt. We hypothesize that there are a small number of facets that collectively define a task, which can be discovered by inspecting a representative set of examples from the data distribution. Consider the task of tagging hate speech. Identifying intent to harm, recognizing general aggression versus targeted hate, and differentiating hate from other objectionable content are some facets of the task. We ask whether there is submodularity structure in adding different facets (text sections) to a prompt, compared to adding in-context examples. The answer to this question depends on multiple factors: (1) what are the contents of the sections and how are they generated?; (2) how does $f(\mathbf{x})$ behave w.r.t. the sections that make up the prompt $\mathbf{x}$?
Figure 2 (b) shows the distribution of marginal utility differences for $f_{\text{facets}}$ (dark blue), i.e., we sample multiple $A,B$ sets of facets (i.e., short sections) that are generated using a Llama2-13B model, and compute $\Delta_{B}-\Delta_{A}$ (defined as above, but the added element $e$ is now a section) using GPT-3.5 as the solver LLM (See Appendix A.2 for experimental details). We see that the diminishing marginal returns property largely holds for $f_{\text{facets}}$, unlike in the case of $f_{\text{fewshot}}$. This suggests that greedy strategies for prompt optimization problem over sections could work — but this crucially relies on how we generate the sections themselves.
The prompt optimization problem can now be posed as one of inferring the facets of a task from a set of training examples. In the following section, we propose a method that iteratively learns the facets of the task as sections of the prompt.
3 UniPrompt: Generating a Prompt that Captures Multiple Facets of a Task
Manual prompt engineering consists broadly of two steps: 1) generating an initial prompt; and 2) identifying errors and editing the prompt. The initial prompt for a task is constructed based on domain knowledge. Typically, the initial prompt contains implicit or explicit sections (e.g., introduction, task description, intended application, edge-cases, challenges, etc.) wherein each section corresponds to a relevant task facet. Given the initial prompt and a solver LLM, a human prompt engineer would look at the incorrect predictions, form hypotheses on which section of the prompt needs to be edited or whether a new section needs to be added, and then propose edits that may fix the errors. Then, the prompt engineer may pick the edit with the highest validation accuracy. These steps are repeated in a loop, until the validation accuracy saturates. A good engineer may also backtrack from an optimization trace and start from some previous prompt, in case the current optimization trajectory yields unsatisfactory results.
We present UniPrompt, a prompt optimization algorithm that aims to mimic the above process. Specifically, the algorithm extracts key concepts or facets relevant to a task and updates prompt sections using them, with the overall goal of increasing validation accuracy. Our analysis from Section 2 indicates that such a directional optimization procedure based on iteratively adding task facets can converge to a good solution, especially for solver LLMs such as GPT-3.5 and GPT-4.
[…]
1) Task Facets Initialization using background knowledge. In the first stage, the algorithm leverages a language model to generate a prompt with multiple facets, i.e., sections focusing on specific task aspects based on the task description. It greedily selects content for each facet from a pool of candidates generated using the language model.
2) Task Facets Refinement using examples. In the second stage, the algorithm refines the prompt using a set of training examples and an auxiliary LLM. Given a candidate prompt, it collects examples with prediction errors by the solver LLM, extracts key concepts or facets required to fix those errors, and makes appropriate edits. An edit may create a new section, modify an existing section, or delete an existing section. Additionally, we maintain a beam of best performing prompts to backtrack and restart from a previous prompt if the current optimization path is unsatisfactory.
Similar to UniPrompt, previous methods [12, 22] propose using an auxiliary LLM to provide feedback on incorrect predictions and propose edits to a given prompt. In our initial experiments, however, we find that collecting feedback over individual examples or randomly sampled batches leads to memorization of individual examples (see Figure 1) rather than recognition of facets that are important to the task. To avoid overfitting to the train examples and to extract generalizable concepts, we make two contributions. First, we follow a two-tier setup of synthesizing feedback for a batch of training examples. We break up a batch into mini-batches, collect feedback on each of the mini-batches and then use a separate prompt to aggregate the different feedback texts into a generalizable concept. Second, to increase chances that a mini-batch corresponds to a coherent facet, we cluster the training data beforehand and ensure that each mini-batch consists of examples from the same cluster. We provide details of the algorithm below (see Algorithm 1).
Notation: We denote the training set with $D_{t}$ where each example is a question-answer pair of the form $\{q_{i},a_{i}\}$. The validation set is denoted as $D_{v}$. We have $N$ training examples and $K$ validation examples. Input to the algorithm is the solver LLM $f_{LLM}$, train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$. In addition, we assume access to an “expert” auxiliary LLM such as GPT-4.
Input: Train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$
Output: Optimized prompt $P^{*}$ for the given task
1 Cluster train set, initialize history and validation accuracy arrays: $C\leftarrow$ cluster($D_{t}$), $H\leftarrow\{\}$, $V\leftarrow$ [];
2 Initialize a beam of size 2 with the initial prompt: $p_{1}\leftarrow p_{0}$ and $p_{2}\leftarrow p_{0}$ for each cluster c in C do
3       $B\leftarrow$ batches($C$);
4       for each batch $b\in B$ do
5             $M\leftarrow$ mini-batches($B$)
6             $F\leftarrow[]$
7             for each mini-batch $m\in M$ do
8                   Evaluate the best prompt on mini-batch: $a_{m}\leftarrow$ LLM($m,p_{1}$)
9                   Get feedback from the expert given history of mini-batch, accuracy and task description: $f\leftarrow$ Feedback($T$, $a_{m}$, $H[m]$)
10                   $F$.insert($f$)
11
12            Combine feedbacks over a batch: $F_{b}\leftarrow$ Combine($F$)
13             Apply feedback to get updated prompts; $q_{1}\leftarrow$ apply($F_{b}$, $p_{1}$); $q_{2}\leftarrow$ apply($F_{b}$, $p_{2}$)
14             Update the beam: if not($p_{1}$ = $p_{0}$) then
15                   $p_{2}\leftarrow$ second-high-acc([$p_{1}$, $p_{2}$, $q_{1}$, $q_{2}$], b)
16            $p_{1}\leftarrow$ highest-acc([$p_{1}$, $q_{1}$, $q_{2}$], b)
17      Evaluate the best prompt on validation set: $acc_{v}\leftarrow$ evaluate($p_{1}$, $D_{v}$)
18       $V\leftarrow V$.append($acc_{v}$)
19       Stop based on the validation accuracy: $c\leftarrow$ early-stop-criteria($V$)
20       if $c$ then
21             break
22
23return $p_{1}$;
Algorithm 1 UniPrompt
3.1 Task facet initialization using background knowledge
While existing methods for prompt optimization [12, 22] iteratively edit a human-written prompt, our goal is to generate the entire prompt from scratch (specifically, based only on a one line description of the task). Therefore, one option is to initialize the prompt using only the task description. We call this Task Description Initialization and initialize the prompt with a single section titled Introduction containing the input task description.
In the second kind of initialization, we finetune a language model to generate a prompt with multiple diverse sections such as Introduction, Tricks, and Corner Cases; similar to the initial prompt that a human prompt engineer may produce. Here we use the world knowledge embedded in language models that are trained on extensive text data. The language model is finetuned to output the entire section contents given the task description and the section title. Specifically, we train a Llama2-13B model using a synthetic dataset generated using GPT-4. Across a wide range of tasks, we first use use GPT-4 to identify the important sections given the task, section descriptions and a few examples (See Appendix A.6 for the prompt). We then fine-tune a language model $\phi$, that is Llama2-13B, in a supervised setting on a synthetically created dataset of the form $\langle T,P_{i},S_{i},{\bf S^{\prime}}_{\text{gold}}\rangle$ where $T$ denotes the task description, $P_{i}$ is the prompt generated using the fine-tuned model sequentially before section $S_{i}$ (i.e., it would have optimized sections $\{S_{1},S_{2},...S_{i-1}\}$), and ${\bf S^{\prime}}_{\text{gold}}$ is the gold section contents. $\phi$ is finetuned on the objective: $\min_{\phi}[-\log(p_{\phi}({\bf S^{\prime}}_{\text{gold}}|T,P_{i},D_{t},S_{i}))]$. We then sample 10 candidates for each section using the fine-tuned model and evaluate each section added incrementally against $D_{v}$ and select the best performing candidate for each section. More details on data collection and training of SLM are given in the Appendix A.4. Examples of one line task description and the Llama-initialized prompt for different tasks are in Appendix A.3.
3.2 Task facet refinement using examples
Extracting task-relevant concepts from a set of examples to refine a prompt is a complex problem comprising multiple steps. Given a set of incorrect predictions, one needs to analyze what went wrong in each prediction, form hypotheses, aggregate the hypotheses to identify specific concepts that are relevant for the task, and then for each concept, one needs to attribute which facet/section of the current prompt needs to be edited to incorporate the concept. These operations are highly model-specific and are difficult to execute reliably. Therefore, for this stage, we exclusively rely on an “expert” auxiliary LLM such as GPT-4.
The whole process is executed through a series of prompts to the expert LLM. First, we prompt the expert LLM to diagnose mistakes (feedback) in each example given the answer and chain-of-thought reasoning produced by the solver LLM. Subsequently, we use this feedback to generate precise edits for the prompt that may fix the error. These individual edits are then aggregated over a mini-batch and fed back into the same LLM, which then identifies a few major edits to be applied to the current prompt. To aid in identifying major edits that correspond to generalizable facets, we propose to cluster the examples as a preprocessing step and create clustered batches, such that each cluster shares some common facet of the task.
3.2.1 Preprocessing step: Clustering examples to facilitate facet identification
We explore two approaches for clustering: topic based clustering, and feedback based clustering.
Topic-Based Clustering. Given a set of examples, we identify $l$ topics that would span the entire train set. This type of clustering is motivated by the observation that solver LLM may make similar mistakes on examples from the same topic. Hence, for such examples, a common edit to the prompt could improve predictions for all the examples. To obtain the clusters, the expert LLM is prompted (for prompt see Appendix A.7) to provide a broad sub-topic $t_{i}$ for each question. Then the resultant list of sub-topics $\{t_{1},t_{2},\dots,t_{N}\}$ is again clustered into $k$ topics $\{t^{\prime}_{1},t^{\prime}_{2},\dots,t^{\prime}_{l}\}$ by prompting the expert LLM. Based on this clustering, each example ${q_{i},a_{i}}$ is assigned a cluster $t^{\prime}_{j}$ corresponding to $t_{i}$.
Feedback-Based Clustering. Another way to find examples that share similar task facets is the feedback they receive based on the initial prompt’s predictions. For instance, for a physics-based task, if two examples from different topics obtain the same feedback from the expert LLM to edit the “Rules” section of the prompt and include the statement, “Draw all forces on each body before writing the equations", then we argue that such examples can be clustered. This type of clustering makes the broad edit identification step easier. To obtain the clusters, we first evaluate all training examples against the initial prompt $p_{0}$ and store the feedback $f_{i}$ from the expert LLM, corresponding to each incorrectly answered example ${q_{i},a_{i}}$ (all the correctly answered questions form one cluster). We then prompt the expert LLM to cluster these feedbacks $\{f_{1},f_{2},\dots,f_{N}\}$ into $l$ clusters $\{f^{\prime}_{1},f^{\prime}_{2},\dots,f^{\prime}_{l}\}$ (see Appendix A.8). For each cluster $f^{\prime}_{j}$, we create a batch of examples ${q_{i},a_{i}}$ corresponding to each $f_{i}$ in $f^{\prime}_{j}$.
3.2.2 Generating prompt edits that generalize to multiple examples
Two-tier Feedback. To encourage generalizable feedback from the expert LLM that applies to more than one example, we obtain feedback at two levels: mini-batch and batch. Given a batch (created using clustering discussed above), we break it up into mini-batches. For each mini-batch $m$, we consider the questions for which the solver LM answered incorrectly with the current prompt. For these questions, we construct a prompt consisting of the questions, the chain-of-thought produced by the solver LLM, their incorrect predicted answers and the ground-truth answers. We ask the expert to provide one feedback for the mini batch (prompt is provided in Appendix A.9). In the prompt, we specify the following types of edits that the expert LLM can suggest: add a section or subsection, delete a section or subsection, and edit a section or subsection.
Given the different feedback texts for mini-batches within a batch $b$, the next step is use another call to the expert LLM to summarize the given set of edits and provide a single summarized section update. This combination ensures some degree of smoothness at every update which helps stabilize training. To make sure that the expert is able to generate generalizable edits, we additionally provide a random set of incorrect examples that are not in the current batch and ask it to suggest an edit based on the existing edits that can correct the errors. As before, the class of edits allowed is the same.
History for effective exploration: To ensure comprehensive and non-repetitive exploration of prompts, we also provide the batch-level history of edits in the mini-batch-level prompt. Specifically, we provide the expert with the history of edits $H[b]$ [12, 31] for the batch $b$ and the corresponding accuracy difference that each edit caused over the previous edit for the given batch. History is presented in the format $\{e_{i},acc_{i}-acc_{i-1}\}$ where $e_{i}$ is the edit proposed at the $i^{th}$ update and $acc_{i}$ is the accuracy of the $i^{th}$ updated prompt (See Appendix A.9 for the full prompt).
3.2.3 Editing the prompt
Once the final set is edits is received for a batch, we use the expert LLM to apply edits to the current prompt (see Appendix A.10 for the prompt). An edit is accepted only if it increases the validation accuracy compared to the current prompt. We call this approach “Greedy”. Alternatively, we maintain a beam of $l$ best performing prompts based on validation accuracy, apply the edit to all the $l$ prompts, and update the beam to retain the top $l$ performing prompts. We call this method “Beam”. To avoid overfitting on the train examples (or keep adding unnecessary information to the prompt), we employ early stopping in the optimization process (More details given in Section 4).
3.3 Efficiency considerations: Computational Complexity
The computational complexity of UniPrompt is dominated by the training phase, and in particular, by the queries made to the solver and the expert LLMs; at inference time, we need a single query to the solver LLM using the learnt prompt and the given test example. This is particularly well-suited for settings such as recommendation systems, where inference time efficiency is critical. In contrast, methods such as MedPrompt [21] require multiple calls at test time for a single test example. Let us now consider the compute complexity in terms of the number of expert or solver LLM calls made per epoch, stage-wise.
Clustering: First, we evaluate all the training examples using the current prompt. Second, for every wrongly predicted example, we obtain feedback from the expert LLM. Third, for the given set of feedbacks, we use a single call to cluster it into $l$ clusters. Each of the above steps incurs $O(N)$ queries, so the total query complexity of the clustering stage is $O(N)$. Finally, for each example, i.e., (question, answer) pair, we simply map it to the $l$ clusters (no LLM calls). This is a one-time cost.
Mini-batch feedback and Batch-level aggregation: At a given epoch, we evaluate every question in the mini-batch using the current prompt and the solver LLM ($N$ queries overall). Next, we obtain one feedback over all the wrong questions in the mini-batch $m$ ($N/|m|$ queries). We use one call to aggregate these feedbacks. For prompt selection, we evaluate $2k$ prompts on the batch $b$, where $k$ is the beam size ($O(2k|b|)$ queries per batch). Hence overall query complexity is $N+N/|m|+2kN+1$ or $O(kN)$. In practice, $k$ is a small number (2 in our experiments).
Assuming LLM throughput of 0.5 queries per second, a training + validation set of 300 examples, $l=10$ clusters, beam size $k=2$ and 20 epochs, it takes under 7 hours to train.
4 Experiments
Datasets We perform evaluation on four standard datasets : (1) Ethos [20], (2) ARC [4], (3) MedQA [13], and (4) GSM8K [5]. Ethos, ARC, and MedQA contain multiple choice questions, and GSM8K contains questions with integer answers.
Implementation details We set the initial prompt $p_{0}$ for each task as the one line task description and use 200 data points $(q_{i},a_{i})$ as the train set. We use GPT-3.5-Turbo as the solver model, that is used in a chain-of-thought setting. For Feedback and Combine, we use GPT-4 as the expert. We maintain a beam size of 2. Mini-batch sizes (and batch sizes) are constrained by the context length of GPT-4. We find that mini-batch sizes 3 to 5 and batch sizes 5 to 7 work the best for our datasets. The temperature of the LLMs for our method is set to 0 for reproducibility of results. We employ early stopping at batch-level in UniPrompt. We find that a $patience$ value of 5 works best for the given tasks.
Baselines We compare UniPrompt with the following techniques: (1) Task Description: prompt is the one line task description that we use to initialize UniPrompt; (2) Chain-Of-Thought (or CoT) prompting [14]; (3) Expert Prompt: the prompt optimized by humans taken from prior works [21]; (4) OPRO [31], that uses LLMs for discrete optimization over text prompts; (5) ProTeGi [22] that proposes textual gradients and selects edits to prompts using bandit techniques; (6) Evoke [12] that uses two instances of LLM, one that scores the current prompt, and the other that edits the prompt; (7) EvoPrompt [7] that uses genetic algorithms to search through the space of prompts; and (8) MedPrompt [21] that effectively combines multiple prompting strategies such as kNN-based few-shot selection, CoT reasoning, and ensembling.
Metrics For each dataset, we perform 5-fold validation (we split the train set into 5 folds) and report the final test accuracy as the average accuracy of the 5 runs on the test sets.
4.1 How does UniPrompt compare to existing methods in zero-shot setting
In the zero-shot setting, we do not include labeled examples in the prompt for any of the compared methods. We report results for four versions of our method in Table 1; they differ in the initial prompt—task description, expert prompt, or the prompt generated by the fine-tuned Llama model (discussed in Section 3.1)—and the combining strategy (from Section 3.2.3)—beam search vs greedy.
UniPrompt variants significantly outperform the baselines including CoT and the state-of-the-art prompt optimization techniques like ProTeGi that crucially leverage LLMs for performing iterative prompt edits. We achieve maximum gains on the Ethos dataset with a $18.2\%$ increase in accuracy over the expert prompt. Further, we see accuracy increases of $4.0\%$ on MedQA, $3.5\%$ on GSM8k, and $7.6\%$ on ARC-Challange datasets.
An example of evolution of prompts using our algorithm is given in Appendix4. It starts with a simple description of task and adds important details like differentiating between hate speech and rudeness. In contrast, ProTeGi [22] yields a rather terse prompt on the same dataset: “Does the following text contain language that targets a group of people based on their religion, gender, or other personal characteristics?”. This prompt does not provide various facets of the task, the way our method infers or an expert would typically write.
Table 1: Test accuracies for the compared methods with GPT-3.5-Turbo as the solver model in the zero-shot setting (best in bold; second best underlined). The last four rows correspond to our proposed method.
Method Ethos ARC MedQA GSM8K
Task Description 76.8 79.7 52.7 59.4
Expert Prompt 74.1 78.4 53.1 78.9
Llama Prompt (Section 3.1) 74.0 89.7 52.6 79.5
CoT 72.0 79.4 50.3 76.3
OPRO 65.4 79.1 53.3 77.1
ProTeGi 76.0 78.8 52.9 77.3
Evoke 63.5 89.0 52.8 81.0
EvoPrompt 81.6 89.9 50.3 81.4
UniPrompt (Init = Expert Prompt) + Beam 84.0 86.0 52.3 82.4
UniPrompt (Init = Task Description) + Beam 92.3 86.0 57.1 82.4
UniPrompt (Init = Llama Prompt) + Beam 92.0 90.5 55.5 81.5
UniPrompt (Init = Task Description) + Greedy 93.7 90.5 55.5 82.3
The training curves in Figure 3 show that our method initially performs edits on the prompt that simultaneously increase the train as well as the validation accuracy. After about 10 or 15 iterations (each batch update is an iteration), validation accuracy decreases while train accuracy continues increasing, indicating overfitting; which we overcome by using early stopping.
4.2 How does UniPrompt perform compared to SOTA in the few-shot setting?
We now compare the performance of UniPrompt, in the few-shot setting with GPT-4 as the solver model, to the SOTA prompting technique MedPrompt [21] that employs three key components: (1) few-shot prompting, where they dynamically select five relevant examples using k-nearest neighbors (kNN) clustering in the embedding space; (2) Chain-of-Thought (CoT) reasoning on the selected examples; and (3) self-consistency and ensembling with option shuffling. From Table 2, we observe that UniPrompt (first col), which requires only one call at inference time, performs almost as well as MedPrompt (last col), which requires five calls, on the MedQA, PubMedQA, and MedMCQA datasets. As we incrementally add components such as kNN few-shot, CoT, and Ensemble to our prompt, we see a significant increase in accuracy of $4.35\%$ on average across all datasets.
4.3 Results on a Real-world Task: Search Query Intent
We consider the task of inferring if two search queries share identical intent or not. This is an important task arising in search and recommendation pipelines, and is tackled today using LLMs. It is challenging to tell if two queries share the same intent, as it requires domain knowledge (e.g., brands and product categories), depends on geographical biases (e.g., “cricket” and “cricket game” are likely to mean the same in UK, but unlikely in the US), and even on cultural connotation. So, examples are crucial for understanding the task and engineering a prompt that generalizes well.
We sample real user queries from a proprietary application, rewrite them using ML models, and ask expert judges to label the query-pairs as identical or otherwise based on prescribed guidelines. We use a set of 200 examples as training data, and an additional 50 examples as validation set, to learn a prompt using UniPrompt, starting from the one-line description: Tell if query A and query B have same intent or not. The dataset is heavily biased towards positive samples, so the metric of success is improvement in accuracy, over the best manually-engineered prompt, on the positive and negative classes individually. For testing, we use a separate labelled set of 2527 examples from two geographies — one where the training data was sampled from, and the other unseen.
The prompt obtained using UniPrompt improves over the best manual prompt by 5.77% on the negative (rare) class, by 0.23% on the positive class, and by 1.86% overall on the test set. The learnt prompt captures the following facets of the task: (1) recognizing variations in names and abbreviations, and how they do not necessarily change the context; (2) recognizing the specificity of brands, and how even minor variations do change the context; and (3) recognizing the specificity of terms in queries, and how lack of specific terms can indicate departure of intent.
4.4 Do diverse task facets organized as sections really help?
We now revisit our hypothesis and motivation in Section 2.2. We want to empirically validate if all the diverse task facets that UniPrompt learns indeed contribute to the performance gains that we observe in Table 1. We consider two ablations:
1) We successively remove each facet (i.e., sections) in the learnt prompt for the task and report the performances of the prompts with fewer facets. In Figure 3, for the Ethos dataset, we see that almost every additional facet contributes to non-trivial gains in accuracy.
2) Could we have captured the information differently and retained the performance? We do a simple experiment – we summarize all the facets (i.e., learnt prompt) and evaluate the resulting prompt. In Figure 3 (right) (green line), we see that the summarized prompt has a significant accuracy drop.
4.5 Ablations: Impact of Clustering, Inclusion of History, and Greedy Update
Here we study effect of our design choices, namely clustering (Section 3.2), inclusion of history of past prompts (Section 3.2), and the greedy update rule (Section 3.2.3). The results are shown in Table 3. We see that clustering as well as edit history components are critical for performance of UniPrompt in all the datasets. We see a major drop of $14.8\%$ in accuracy in the Ethos dataset when clustering is removed, and a $4.3\%$ drop when history component is removed. In all the datasets except GSM8K, we find clustering is more important than history. This can attributed to variability of question types being less in GSM8K than in others: all questions are grade-8 arithmetic questions in GSM8K.
We also find that the greedy update rule proves to be superior or competitive compared to beam search in relatively easier datasets — where even less exploration produces good results, greedy proves to be a more effective update rule. On the other hand, in more complex datasets like MedQA, greedy appears to be a bad strategy. We also see that clustering examples based on feedback (“Fb Clustering”) is a better strategy than clustering based on topics, except for the Ethos dataset.
5 Related Work
Here, we highlight relevant work that are not addressed in the manuscript so far. Deng et al. [6] present a discrete prompt optimization method, RLPrompt, using reinforcement learning, where a policy network learns to generate effective prompts through reward-based training, with an emphasis on enhancing training efficiency through effective reward stabilization techniques. A drawback of such automatic prompt optimization approaches [22, 36, 6, 31] is that the prompts generated tend to be short, often comprising only one or two sentences, which may not fully encapsulate the complexity of the task at hand.
Prior research [28, 27] has highlighted the significance of specific sections within prompts. For instance, a well-designed prompt typically incorporates elements such as a system prompt and a chain-of-thought section. However, existing methods do not specifically target the optimization of individual sections and their respective contents within the prompts. Hsieh et al. [11] investigate the use of greedy and genetic algorithms to edit lengthy prompts. Their method focuses on paraphrasing one line at a time starting from an existing prompt, compared to our goal of learning facets of a task from scratch, thereby compromising generalization accuracy. Another orthogonal line of work that ha been explored recently for prompt optimization is the algorithmic selection of in-context examples [19, 9, 30, 25, 26].
6 Limitations and Future Work
We presented a method inspired by the human prompt engineering process to generate complex prompts from scratch that include different facets of a task. Our algorithm provides significant improvements over baseline prompt generation methods on multiple standard datasets. However, in our evaluation, we considered only classification tasks and only GPT-4 as the expert LLM. Generalizing to generative tasks, as well as investigating the impact of the expert LLM choice are some future directions.
Appendix A Appendix
Table 2: Comparison of UniPrompt (“Ours”) with MedPrompt, with GPT-4 as the solver model.
Ours Ours + KNN Ours + KNN + COT Ours + KNN+COT+Ensemble MedPrompt
MedQA 80.9 81.0 83.9 87.0 80.6
PubMedQA 70.3 72.2 74.7 75.6 71.2
MedMCQA 79.2 81.4 82.6 84.5 79.1
MMLU MG 78.0 94.0 96.0 99.0 98.0
Figure 3: Training curves for MedQA (left) and ARC (middle) datasets when UniPrompt is initialized with (published) state-of-the-art prompts; (right) ablation of facets on Ethos (Section 4.4).
Table 3: Ablation of design choices in UniPrompt with GPT-3.5-Turbo as the solver model.
Ethos ARC MedQA GSM8K
UniPrompt -- History 88.0 84.6 55.3 80.8
UniPrompt -- Clustering 77.5 82.0 54.1 81.5
UniPrompt 92.3 86.0 57.1 82.4
UniPrompt $+$ Greedy 93.7 90.5 55.5 82.3
UniPrompt $+$ Fb Clustering 87.2 91.2 58.3 82.5
A.1 Details on estimation of Lipschitz constant $L$
TO calculate the Lipschitz constant for a given LLM and task, we take a human written promp and generate it’s paraphrases using GPT-4. We prompt GPT-4 with the following text: “You are given a sentence, you have to generate 30 paraphrases of the sentence, make sure that the cor content of each paraphrase is same, you can use add, subtract or change words". These paraphrases are then evaluated on the validation set $D_{v}$. For a measure of distance between two prompts, we take the cosine similarity between the embeddings of two prompts. We use text-ada-002 for generating the text embeddings for prompts.
A.2 Details on marginal utility difference calculation for prompts
For a given LLM and task, to study the marginal utility of different facets for a prompt, we first sample a few facets (generated using the fine-tuned model) that individually perform better than the chain-of-thought prompt as a baseline. Given these set of facets, we sample three facets $f_{1},f_{2},f_{3}$ and generate four prompts: $A:f_{1},B:f_{2},A+\Delta:f_{1}+f_{3},B+\Delta:f_{2}+f_{3}$. We define $\Delta_{A}$ as the accuracy difference between prompt $A$ and prompt $A+\Delta:f_{A}+f_{A+\Delta}$. Similarly, $\Delta_{B}=f_{B}+f_{B+\Delta}$
A.3 Prompt Initialization
One line task descriptions:
  1. 1.
    Ethos: In this task, you have to determine whether a given text is hate speech or not.
  2. 2.
    ARC: You have to solve the following science question.
  3. 3.
    GSM8K: In this task, you are given a math question. You have to solve the question.
  4. 4.
    MedQA: In this task, you are given a medical question. You have to solve the question.
An example of sectioned initialization prompt generated using finetuned Llama Model
Introduction:
Assume the role of a science expert and answer the given
question by selecting one of the options A, B, C or D.
         1. Understand and solve science questions by selecting
         the best answer from a given list of options.
         2. Identify the logic behind the choices provided and
         make an informed decision.
         3. Use contextual clues to choose the most accurate
         answer.
         4. Be aware of the differences between science and
         everyday language.
Task Description:
          Scientific inquiry: Science is the systematic study of
          the structure and behavior of the physical and natural
          world through observation and experiment. The
          scientific method is a process for acquiring knowledge
          that has been improved upon since its inception in the
          17th century. It involves making observations,
          formulating hypotheses as to their causes, and
          experimenting with them to support or refute the
          hypotheses.
Real-life Application:
1. Assisting Students in Science Classes:
    In the context of science education, the ability to solve
    science questions can help students to better understand and
    internalize the concepts. By familiarizing themselves with
    the basic principles of science, students can develop a
    stronger foundation of knowledge.
2. Improving Scientific Literacy:
    Scientific literacy is a critical skill in today’s world,
    where scientific knowledge is increasingly important. By
    solving science questions, individuals can improve their
    understanding of scientific concepts and be more informed
    about scientific developments.
3. Scientific Questions:
    In daily life, there are many questions that require
    scientific knowledge to answer. For example, understanding
    the science behind certain phenomena, such as why a magnet
    sticks to a refrigerator door, can help us in our day-to-day
    life.
4. Increased Awareness:
    By answering scientific questions, we can develop a deeper
    understanding of the world around us and increase our
    awareness of scientific phenomena. This can help us in our
    daily lives and make us more knowledgeable individuals.
Background Knowledge:
    1. Understanding of the basic concepts of science and
    physics, such as the difference between heat, temperature
    and friction.
    2. Basic knowledge of the different types of skin surfaces,
    such as dry, wet, rough, smooth, etc.
    3. Familiarity with the different types of magnets and their
    properties.
    4. Understanding of the different factors that affect the
    adhesion of magnets to different surfaces.
    5. Knowledge of the different types of sedimentary rocks and
    their properties.
Challenges:
1. Ambiguity in the question:
    The question might be ambiguous in nature, and it can be
    difficult to understand the exact meaning of the question.
    In such cases, it is important to read the question
    carefully and identify the key concepts or keywords. This
    can help in arriving at the correct answer.
2. Scientific terms or concepts:
    The question might contain scientific terms or concepts that
    are unfamiliar to the user. In such cases, it is important
    to understand the meaning of these terms or concepts and
    their relationship with the question.
3. Difficulty in understanding the question:
    Sometimes, the question might be complex or abstract, making
    it difficult to understand or interpret.
4. Misleading statements or information:
    The question might contain misleading or false information,
    making it difficult to determine the correct answer.
5. Contradiction:
    The answer can be in conflict with well-known scientific
    facts or principles. In such cases, it is important to make
    a careful analysis of the evidence and choose the answer
    that is most consistent with the available
Simplification:
1. Identify the key elements in the question:
    Ask yourself, "What is the main question in the question?"
    Identify the key elements and focus on them to solve the
    problem.
2. Understand the context:
    Understand the context of the question and the background
    knowledge you need to answer it.
3. Identify the answer choice:
    Identify the answer choice that best fits the context and
    background knowledge.
4. Eliminate the distractors:
    Eliminate the distractors that don’t fit
Tricks:
1. Read the question carefully: Understand the question and its
context. This will help in understanding the information and
concepts needed to solve the question.
2. Identify the key concepts: Identify the key concepts and
keywords in the question. This will help in understanding the
main idea and focus on the relevant information.
3. Understand the question structure: Understand the structure
of the question. This will help in identifying the appropriate
answer option and avoiding distractions.
4. Look for clues: Look for clues in the question and the
answer options
A.4 SLM Training Details
To induce the ability of structured prompt generation in a smaller language model, we curate a section-wise dataset of around 12,000 task-prompt pairs. The tasks for training dataset creation were taken from tasksource library [24] that contains around five hundred classification tasks. We extract the task description from tasksource-instruct, which contains tasksource dataset recasted with instructions. For instance, the task description for BIG-bench Entailed Polarity task is, "Given a fact, answer the following question with a yes or a no". The dataset provides diverse tasks and their short description, but not the human-generated prompts for each task. To approximate human-generated prompts, we use GPT-4 as a teacher model.
By prompting GPT-4 with the task description and section description, we ask it to generate the contents of the section. To ensure that the generated section-wise prompts are concise and relevant, we prompt GPT-4 to not generate more than five lines of content for each section. We use LLAMA2-13B model, which we finetune using LoRA adapters as the auxiliary LM that generates sections.
A.5 Prompt to Llama2-13B for fine-tuning
### Instruction:
You are a prompt engineer, you have to write a structured prompt.
For the given task description, examples and section description,
write the contents of the section that align with section description.
### Task Description:
{data_point[’task_description’]}
### Section Description:
{data_point[’section’]}:{section_descriptions[data_point[’section’]]}
### Response:
{data_point[’prompt’]}
A.5.1 Examples of training dataset
Task: super glue boolq
Task Description: With no explanation, label the following with
either False or True.
Section: description
Prompt:  As a prompt engineer, your task is to create a prompt
that accurately labels statements as either True or False
without providing any explanations. The statements may involve
general knowledge, trivia, or facts about events, places,
people, or objects. Ensure you have a strong understanding of
the subject matter and write a concise and specific task
description related to the given statements.
Task: glue qnli
Task Description: With no explanation, label A to B with either
entailment or not entailment
Section: background
Prompt: 1. Entailment means that the information in statement B
can be inferred directly from statement A.
2. Not entailment means that the information in statement B
cannot be inferred directly from statement A or is unrelated.
3. Understanding the context and relationship between the two
statements is crucial for accurate classification.
Task: bigbench undo permutation
Task Description: In the following sentences with shuffled
words, find the correct order of word swaps to unscramble the
sentence.
Section: tricks
Prompt: 1. Identify the key words or phrases in the task to
understand the context of the sentence. Look for nouns, verbs,
and adjectives that seem related or could logically fit together.
2. Start by solving the problem step by step and focus on one
swap at a time. Breaking the problem into smaller sub-problems
will make it easier to manage.
3. To make the task more manageable, first focus on swapping the
words that are clearly out of place, such as words that should
be at the beginning or end of the sentence.
A.6 Prompt for identifying important facets
you are given a task, along with it’s description, some examples
of how to solve the task and section descriptions.
What do you think would be the most important sections to
include for the given task.
## Task
{task}
## Task Descirption
{tas_description}
## Examples
{Examples_string}
## Section Descriptions
{sections}
A.7 Clustering Type 1
You are given a science question, you need to tell which broad
topic is this question from.
Question: {train_questions_new[ij]}
Answer: {answer}
Give your answer as a single word, between <Answer></Answer>
tags like: <Answer>Thermodynmics</Answer> or
<Answer>Botany</Answer>.
Subtopic:
A.8 Clustering Type 2
You are given a set of feedbacks, you need to cluster them into
five groups based on similarity, and then provide a summary of
[…]
{feedback}
provide each cluster explnation within the following tags:
<Cluster></Cluster>
You are given a feedback and a set of clusters, you need to tell
which cluster this feedback belongs to.
[…]
The feedback is: {feedback}
give your final answer as the number of the correct cluster
between <Answer></Answer> tags like: <Answer>1</Answer>.’’’
A.9 Feedback Prompts
Feedback over mini-batch
You are a teacher and you have to give feedback to your students on their answers.
You are teaching how to solve math problems to your students.
You are given a question, it’s true answer and answer given by student.
You are also given the explanations written by your students while solving the questions.
The questions are answered wrong by the students.
You have to tell why is the solution wrong and what information
is can be added to the in the Background Knowledge part that
would have helped the student to write better explanations.
## IMPORTANT: You are also given a history of changes you made
to the background knowledge part and the change in student’s
accuracy after making the change. You have to use this history
to make your feedback.
Be explicit and tell the exact information that can be added
without further modification / addition.
### IMPORTANT: Give feedback in form of instructions like  add a
section, add a subsection, set the content of a section, set the
content of a subsection, delete a section or delete a subsection
in the background knowledge part.
Give very granular feedbacks, like if the student has made a
mistake in the calculation, then tell what is the mistake in the
calculation and how to correct it, if the student has made a
mistake in the concept, then tell what is the mistake in the
concept and how to correct it.
## Background Knowledge
     {current_prompt}
## History
    {history_string}
Now, it is your turn to give feedbacks to the students.
You can only provide a one line feedback.
Feedback over batch
You are given a set of feedbacks for some problems. The set
feedbacks for each problem separated by =========== symbol.
You have to summarize the feedbacks into a final feedback.
You are also given a set of wrong questions. You need to tell
which edit can be applied to aid the student in solving the wrong question.
To achieve your task, try to follow the following steps;
1. Identify the general problem that is being solved by all the
feedbacks.
2. Once you have identified the problem, try to make a new
feedback that covers most of the
feedbacks given.
Let’s say the problem in the first feedback is the absence of
methods to solve linear equation and in the second feedback it
is the method to inverse a matrix.
You know that both of these problems can be caused by adding how
to solve convert a matrix  into row rediced echolon form. So,
add that.
3. Try and validate your feedback. Once, you have a feedback try
to see if it covers every
feedback, if it does not cover any feedback, add that to your
new feedback.
4. See the wrong questions and try to identify what is the
problem in the question.
If the problem is not covered by your feedback, add that to your
feedback.
5. You can add specifics like examples, definitions etc make
sure that the feedback is enough to be directly added without
any modification.
You may use the following function templates-
add_section(sectioname)
add_subsection(section_name, subsection_name)
set_section_content(section_name, new_content)
set_subsection_content(section_name, subsection_name, new_content)
delete_section(section_name)
delete_subsection(section_name, subsection_name)
Your summary cannot include more than four functions. Make sure that the content is useful,
not just a very general statement. Something specific.
Instructions:
{edits}
Wrong Questions:
{wrong_examples_string}
Summary:
A.10 Editing Prompt
You are given an input prompt and a feedback, you have to
incorporate the feedback into the input prompt and output the
final prompt.
An example of the task is given below
### Input Prompt
Introduction: In this task you have to answer the given question.
### Feedback
The background knowledge is incomplete, it does not include what
are the factors that affect the water usage and how many water
sources are there.
[…]
climate, economic development, and availability of water
sources. There are two sources of water, surface water and
groundwater.)
### Final Prompt
Introduction: In this task you have to answer the given question.
Background Knowledge: water usage depends on the population,
climate, economic development, and availability of water
sources. There are two sources of water, surface water and
groundwater.
Only output the final prompt nothing else.
### INPUT PROMPT
{current_prompt}
### FEEDBACK
{edits}
### FINAL PROMPT
A.11 Example of prompt evolution using our method
See example in Figure 4.
Figure 4: Evolution of prompts through iterations of UniPrompt on the Ethos dataset. Starting from a simple one-line prompt having an accuracy of $82\%$, UniPrompt adds background knowledge, corner cases, and additional sub-sections yielding a prompt with accuracy $88\%$. After further iterations, our algorithm converges to a detailed, human-like longform prompt that achieves accuracy of $92\%$.
A.12 Comparision of our method with existing methods
See Figure 5.
Human Prompt
Let’s differentiate using step by step reasoning like a medical expert.
Our Prompt
Introduction: In this task, you are given a medical question. You have to solve the question.
Description: To solve medical questions effectively, it is important to understand various medical conditions, their progression, and associated clinical features.
Background Knowledge: Differential Diagnosis of Subcutaneous Nodules:
When evaluating subcutaneous nodules, consider mobility, consistency, and skin adherence. Epidermoid cysts are firm, non-tender, and the skin cannot be pinched over them. Lipomas are soft, mobile, and have pinchable skin.
Corner Cases: Antiretroviral Therapy Complications:
Doctor should be aware of the common side effects of antiretroviral drugs, with specific attention to the association between didanosine and pancreatitis, and the recommended management strategies, such as replacing didanosine with lamivudine.
Figure 5: Comparison of human-written Prompt and prompt produced by UniPrompt on MedQA dataset.
OPRO optimized prompt
Start by dissecting the problem to highlight important numbers and their relations. Decide on the necessary mathematical operations like addition, subtraction, multiplication, or division, required for resolution. Implement these operations, keeping in mind any units or conditions. Round off by ensuring your solution fits the context of the problem to ensure accuracy
Our Prompt
Introduction: In this task, you are given a math question. You have to solve the question.
Strategies for Word Problems:
1. Understanding Word Problems: When solving word problems, it is crucial to read each sentence carefully and comprehend the time periods and quantities involved. Avoid incorrect multiplication or addition by paying close attention to whether a quantity remains constant over a period or changes. If a quantity is consistent, it does not need to be multiplied by the number of days or weeks unless the problem specifies otherwise.
2. Calculating Averages: To calculate the average of a set of numbers, add all the numbers together and then divide by the number of items. In word problems, ensure you have the correct total before dividing by the number of periods, such as weeks, to find the average for each period.
3. Understanding Past and Future Events in Word Problems: Distinguish between past and future events by identifying the starting and ending points. To calculate the time interval between two events, determine the direction of time from past to future and compute the interval accordingly. This understanding is essential when dealing with problems that ask for the time since a past event or until a future event.
Figure 6: Comparison of prompt produced by the state-of-the-art ORPO [31] and by UniPrompt on the GSM8K dataset.
A.13 Effect of length on performance of prompt
Here we answer the question: How much does only length contribute to UniPrompt’s success?. To answer this, we replace the prompt with in-context examples of the same context length and compare the accuracies in Table 4. We also compare the case where we include only the examples that the solver LLM gives incorrect prediction on, denoted as “Wrong ICL” row in the table. We see that there is a slight increase in accuracy when wrong examples are included in the prompt over randomly including examples. But, overall, UniPrompt performs much better than including in-context examples. This shows that length is not the only factor contributing to UniPrompt’s success.
Table 4: Analysis of the effect of length and contents on the performance of UniPrompt
Ethos ARC GSM8K
UniPrompt 93.7 90.5 82.4
ICL Prompt 63.0 86.7 76.3
Wrong ICL 70.4 87.1 78.2
Summarized Prompt 84.3 85.5 66.0

### v2

@@
 Abstract
-Given a task in the form of a basic description and its training examples, prompt optimization is the problem of synthesizing the given information into a text prompt for a large language model (LLM).
+Given a task in the form of a basic description and its training examples, prompt optimization is the problem of synthesizing the given information into a text prompt for a large language model.
 Humans solve this problem by also considering the different facets that define a task (e.g., counter-examples, explanations, analogies) and including them in the prompt.
@@
 In this work, we view prompt optimization as that of learning multiple facets of a task from a set of training examples.
-We identify and exploit structure in the prompt optimization problem — first, we find that prompts can be broken down into loosely coupled semantic sections that have a relatively independent effect on the prompt’s performance; second, we cluster the input space and use clustered batches so that the optimization procedure can learn the different facets of a task across batches.
-The resulting algorithm, UniPrompt, consists of a generative model to generate initial candidates for each prompt section; and a feedback mechanism that aggregates suggested edits from multiple mini-batches into a conceptual description for the section.
+We exploit structure in the prompt optimization problem and break down a prompt into loosely coupled semantic sections.
+The proposed algorithm, UniPrompt, (1) clusters the input space and uses clustered batches so that each batch likely corresponds to a different facet of the task, and (2) utilizes a feedback mechanism to propose adding, editing or deleting a section, which in turn is aggregated over a batch to capture generalizable facets.
 Empirical evaluation on multiple datasets and a real-world task shows that prompts generated using UniPrompt obtain higher accuracy than human-tuned prompts and those from state-of-the-art methods.
 In particular, our algorithm can generate long, complex prompts that existing methods are unable to generate.
-Code for UniPrompt will be available at https://aka.ms/uniprompt.
+Code for UniPrompt is available at https://aka.ms/uniprompt.
+Task Facet Learning: A Structured Approach To Prompt Optimization
+Gurusha Juneja1,*, Gautam Jajoo2,*, Nagarajan Natarajan3, Hua Li4, Jian Jiao4, Amit Sharma3,† 1UC Santa Barbara, 2BITS Pilani, 3Microsoft Research, 4Microsoft Bing Ads *These authors contributed equally to this work.
+Corresponding author: amshar@microsoft.com
+55footnotetext: To appear in ACL Findings 2025.
 1 Introduction
-Given a task, choosing an input prompt is a key part of optimizing Large Language Model’s (LLM) performance [14, 31].
-Minor changes in prompt can lead to performance gains or losses, necessitating prompt engineering [16] for effective model utilization.
-Typically, manually-developed prompts combine task description with a few in-context examples, along with modifiers like chain-of-thought [14].
+Given a task, choosing an input prompt is a key part of optimizing Large Language Model’s (LLM) performance (Kojima et al., 2024; Yang et al., 2023).
+Minor changes in prompt can lead to performance gains or losses, necessitating prompt engineering (Liu et al., 2023).
+Typically, manually-developed prompts combine task description with a few in-context examples, along with modifiers like chain-of-thought (Kojima et al., 2024).
 For greater accuracy, human prompt engineers spend considerable time to identify errors with a current prompt, consider the different facets of a task (e.g., counter-examples, explanations, analogies) that may fix those errors, and include them in the prompt.
 For instance, for a hate speech classification task, in addition to the definition, it may be helpful to specify the facets that lead to hate speech: the context of conversation, identifying intent, and differentiating hate speech from opinions or closely-related concepts such as vulgarity and profanity.
-To avoid the cumbersome process of manually creating prompts, recent work aims to automate the process of generating natural language prompts that are also interpretable.
+To avoid the above cumbersome manual process, recent work aims to automate the process of generating natural language prompts that are also interpretable.
 Since language tokens are discrete, this leads to a challenging discrete optimization problem with a combinatorial space of possible outputs.
-Techniques for prompt optimization can be divided in two categories: non-directional, e.g., random search [36, 35] and genetic algorithms [31, 7], where the sampling of new input is “random” and does not explicitly aim to reduce error on a train set; and directional, where the sampling of new input depends on some error measure on a representative train sample.
-Recently, more complex methods are proposed in the second category including reinforcement learning [33, 6] and updating prompts using auxiliary LLMs to reduce incorrect predictions over a train sample [12, 22].
-While all these techniques focus on editing, adding, or deleting parts of a given prompt, they are developed with the goal of obtaining a concise description of the task.
-None of them focus on ensuring that multiple facets of a task are added to the prompt.
-In this paper, we study how to optimize prompts to cover diverse, multiple facets of a task and improve overall accuracy.
-To develop a suitable algorithm, we analyze properties of the prompt optimization problem.
-First, we check how sensitive LLMs are to small changes in the input prompt, since directional optimization algorithms may not be as useful if LLMs are highly sensitive.
-Using a notion of probabilistic Lipschitz continuity, we find that LLMs become less sensitive with size.
-Across prompt edits with the same semantic meaning, Llama-13B accuracy fluctuates up to 63% of the minimum accuracy whereas the relative accuracy change in GPT-3.5 and GPT-4 models is 6% and 4% respectively.
-As a result, probabilistic Lipschitz constant for GPT-3.5 and GPT-4 is significantly smaller than that for Llama-13B, indicating that directional optimization methods may be best suited for larger models.
-Second, we study how LLM accuracy changes as more information is added to the prompt.
-Ideally, we would want a diminishing marginal returns (sub-modularity) property so that greedy addition of different facets can lead to a reasonable approximation of the true optimum.
-We find that diminishing returns property largely holds when adding different facets as new sections to the prompt, but it does not hold when directly adding in-context examples.
-Our analysis provides two insights: 1) directional optimization may be feasible for models like GPT-3.5 having a low probabilistic Lipschitz constant; 2) including facets as additional text sections may lead to a well-behaved optimization problem.
-Based on these, we propose an algorithm, UniPrompt, that can generate the full prompt from scratch, given a one line description for a task.
-Specifically, we identify a prompt structure that provides different task facets as sections in the prompt and initializes them with text generated from a language model.
-To develop a high-accuracy prompt, we use an auxiliary LLM to obtain feedback about example predictions with the current prompt and decide to make edits at a section level (add, edit or delete) in each iteration.
-Our key insight is that feedback on a single example or a randomly selected batch of examples does not yield generalizable facet descriptions.
-Instead, we propose generating edits at a minibatch level and aggregating the edits at the batch level to yield a single edit that conveys a broad concept relevant to the task.
-To further increase the chances of facet learning, we cluster the input examples beforehand and use clustered mini-batches (see Figure 1).
-We evaluate UniPrompt on zero-shot (no in-context examples) and few-shot settings on four different tasks.
-In the zero-shot setting, UniPrompt consistently achieves higher accuracy than existing prompt optimization methods.
+Techniques for prompt optimization can be divided in two categories: non-directional, e.g., random search (Zhou et al., 2022; Zhang et al., 2023) and genetic algorithms (Yang et al., 2023; Guo et al., 2023), where the sampling of new input is “random” and does not explicitly aim to reduce error on a train set; and directional, where the sampling of new input depends on some error measure on a representative train sample.
+Recently, more complex methods have been proposed in the second category including RL (Zhang et al., 2022; Deng et al., 2022), updating prompts using feedback from auxiliary LLMs (Hu et al., 2024; Pryzant et al., 2023), and optimizing the input to a small LM that generates the prompt (Lin et al., 2024b; Chen et al., 2024).
+While these techniques focus on editing parts of a given prompt, they are developed with the goal of obtaining a concise description of the task.
+None of these focus on ensuring multiple facets of a task are added to the prompt.
+In this paper, we propose UniPrompt, a prompt optimization method to cover diverse, multiple facets of a task and improve overall accuracy.
+To simulate the manual prompt engineering process, we propose that prompts be constructed from individual sections, where each section may correspond to a different facet that humans may consider for the task.
+Prompt editing proceeds at a section-level: we can add, edit or delete a section from the prompt.
+Similar to Pryzant et al.
+(2023); Hu et al.
+(2024), prompt edits are based on an auxiliary LLM’s feedback about example predictions with the current prompt.
+We contribute two key insights in this feedback-based optimization process.
+First, we find that the feedback on a single example or a randomly selected batch of examples does not yield generalizable facet descriptions.
+Instead, we propose clustering the inputs and creating mini-batches such that each mini-batch is sourced from a single cluster.
+Second, even with clustered batches, the feedback tends to overfit to specific examples or their properties.
+To generate a prompt edit that conveys a generalizable concept relevant to the task, we propose generating edits at a mini-batch level and then aggregating them at the batch level to yield the final edit (Figure 1).
+While the two insights may appear simple, we show that they significantly improve extracting diverse task facets.
+We evaluate UniPrompt on several benchmarks where it consistently achieves higher accuracy than existing prompt optimization methods.
 On Ethos, a hate speech dataset, UniPrompt obtains 94% accuracy whereas the next best method obtains 82%.
-In the few-shot setting, we compare UniPrompt to MedPrompt [21], a state-of-the-art prompt composition method.
+Even though UniPrompt focuses only on the instruction and does not include any in-context examples, we find that its instruction-only accuracy is often higher than methods such as DSPy (Khattab et al., 2024) that optimize both.
+In the few-shot setting, we also compare UniPrompt to MedPrompt (Nori et al., 2023), a state-of-the-art prompt composition method.
 We find that UniPrompt, requiring only one LLM call at inference time, obtains the same accuracy as MedPrompt that requires five calls.
 If we allow multiple calls to UniPrompt, we obtain over 4% accuracy gains.
-We also evaluate UniPrompt on a real-world semantic matching task in a web search engine.
-Compared to the best manual prompt, the prompt generated from UniPrompt leads to over 5% increase in accuracy on the negative class and nearly 2% accuracy increase overall.
-Figure 1: Existing prompt optimization methods (left) versus UniPrompt (right) on the SciQ [29] dataset: [Left] State-of-the-art prompt optimization methods such as ProTeGi  [22] sample from the questions wrongly answered by the current prompt, and use an expert LLM (such as GPT-4) to obtain feedback on the mistakes made; This approach tends to overfit to specific examples, as can been seen from the proposed prompts $p^{\prime}$; leading to limited generalization.
-[Right] In contrast, UniPrompt identifies key task facets that generalize across examples by: (1) clustering examples into groups with similar task facets, and (2) employing a two-tier feedback-based update strategy.
-The resulting prompt updates extract generalizable concepts from the specific examples seen.
-2 Characterizing the Prompt Optimization Problem
-Assume access to black-box solver LLM $f:\mathcal{X}\rightarrow\mathbb{R}$ that takes as input a prompt $\mathbf{x}\in\mathcal{X}$ and outputs the average accuracy on a validation set $D_{v}$.
-Since the set of prompts is combinatorially large, we assume that all prompts can be embedded in a fixed $p$-dimensional space such that distance between two prompts in the vector space correspond to their semantic similarity.
-Thus, each prompt $\mathbf{x}$ can be expressed as a vector in $\mathbb{R}^{p}$.
-The prompt optimization problem can be written as $\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x};D_{v})$.
+Finally, we also evaluate UniPrompt on a real-world semantic matching task in a web search engine.
+Compared to the best manual prompt, the prompt generated from UniPrompt leads to over 5% increase in accuracy on the rare class and nearly 2% accuracy increase overall.
+Figure 1: Existing prompt optimization methods (left) versus UniPrompt (right) on the Ethos dataset: [Left] State-of-the-art prompt optimization methods like ProTeGi  (Pryzant et al., 2023) sample from the questions wrongly answered by the current prompt, and use an expert LLM (e.g., GPT-4) to obtain feedback on the mistakes.
+This approach tends to give very general edits or overfits to specific examples.
+[Right] In contrast, UniPrompt identifies key task facets by: (1) clustering examples with similar task facets, and (2) employing a two-tier feedback-based update strategy.
+The resulting prompt updates extract generalizable concepts from the specific examples.
+2 Related Work
+Here, we highlight relevant work that are not addressed in the manuscript otherwise.
+Deng et al.
+(2022) present a discrete prompt optimization method, RLPrompt, using reinforcement learning, where a policy network learns to generate effective prompts through reward-based training, with an emphasis on enhancing training efficiency through effective reward stabilization techniques.
+A drawback of such automatic prompt optimization approaches (Pryzant et al., 2023; Zhou et al., 2022; Deng et al., 2022; Yang et al., 2023) is that the prompts generated tend to be short, often comprising only one or two sentences, which may not fully encapsulate the complexity of the task at hand.
+Another recent line of work leverages human feedback.
+Automated Prompt Optimization with Human Feedback  (Lin et al., 2024a) optimizes prompts for black-box LLMs using human preference feedback.
+Besides the obvious overhead, it might also introduce potential biases.
+Prior research (Wei et al., 2023, 2024) has highlighted the significance of specific sections within prompts.
+However, existing methods do not specifically target the optimization of individual sections and their respective contents within the prompts.
+Hsieh et al.
+(2023) investigate the use of greedy and genetic algorithms to edit lengthy prompts.
+Their method focuses on paraphrasing one line at a time starting from an existing prompt, compared to our goal of learning facets of a task from scratch.
+Another orthogonal line of work explores algorithmic selection of in-context examples (Min et al., 2022; Gupta et al., 2023; Wu et al., 2023; Srivastava et al., 2024; Sun et al., 2024).
+3 UniPrompt: Capturing Task Facets
 […]
-2.1 Is directional optimization feasible?
-Previous work has shown that LLMs can be brittle to their input: changing the prompt slightly can create a significant difference in performance [37].
-To understand whether optimization algorithms may work, it is important to understand how the surface of $f$ looks.
-If $f$ is brittle everywhere to small changes, then random search over prompts (such as in [36]) may be a competitive method; it is unclear whether optimization can help.
-On the other hand, if the change in $f$ given small changes to the input prompt is bounded, then optimization algorithms can hope to perform significantly better than random search.
-More formally, we ask whether the prompt optimization problem is well-conditioned.
-Typically, conditioning can be determined by the Hessian.
-However, since $f$ is assumed to be black-box, we approximate it by measuring sensitivity, or more specifically, Lipschitz continuity near the optimal solution.
-Based on prior work on defining continuity of neural networks [18], we use a probabilistic notion.
-Definition 1
-Probabilistic Lipschitz Continuity [18].
-Given a probability distribution over inputs $\mathcal{X}$, $r\geq 0$, a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(L,\epsilon)$-probabilistically Lipschitz with constant $L\geq 0$, if
-$\Pr_{\mathbf{x},\mathbf{x}^{\prime}\sim\mathcal{X}}[\operatorname{d}(f(\mathbf{x}),f(\mathbf{x}^{\prime}))\leq L.\operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\ |\ \operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\leq r]\geq 1-\epsilon$ (1)
-where $d$ is a distance measure such as $\ell_{1}$ or $\ell_{2}$ norm.
-Note the focus on small changes in input through the parameter $r$.
-Intuitively, the Lipschitz property bounds the maximum change in $f$ given a small change in input prompt.
-Typically, the lower bound of error for any sequential optimization algorithm over $f$ is directly proportional to the Lipschitz constant $L$ [17].
-Therefore, for faster convergence, it is desirable to have a low $L$, especially near the optimal solution.
-Empirically, we estimate $L$ by sampling task-relevant prompts so that they are close to the optimal solution.
-Then we make small changes to the prompt such that the semantic meaning stays the same and measure the change in $f$.
-We show the change in $f$ per change in input for GPT-4, GPT-3.5, and Llama2-13B models in Figure 2(a) for the Ethos dataset commonly used in prompt optimization [22, 19, 32, 23, 10].
-Assuming $\epsilon=0.05$, probabilistic Lipschitz constant $L$ for GPT-3.5 and GPT-4 is $<1$, whereas it is over three times more for Llama2-13B.
-Thus, as the size of the model increases, the probabilistic Lipschitz constant decreases, indicating that larger models like GPT-3.5 and GPT-4 are more robust to small changes and thus amenable to optimization algorithms.
-For Llama, the absolute change in accuracy ranges between 0.33 to 0.54, indicating a 63% change in accuracy even across prompts with the same semantic meaning.
-In contrast, the relative change in accuracy for GPT-3.5 and GPT-4 is 6% and 4% respectively.
-(See Appendix A.1 for experimental details)
-Figure 2: On the Ethos dataset, (a) estimating (probabilistic) Lipschitz constant of models (Definition 1); (b) diminishing marginal utility (submodularity) of GPT-3.5 on two types of prompting strategies.
-2.2 Can higher-order abstractions of task examples help prompt optimization?
-Consider in-context learning [3], where the prompt $\mathbf{x}$ is composed of a simple task description and a set of labeled examples from the task.
-The optimal prompt should consider all the aspects of the task which can only be discovered by exhaustively looking at the examples.
-But, in practice, we are constrained by (a) availability of a limited number of training examples, (b) prompt size.
-It is beneficial for in-context learning to consider the marginal utility of examples [34, 8].
-For instance, it would help to add examples where the model fails rather than ones where the model already succeeds (see Table 4 in Appendix).
-This suggests that one could use a greedy algorithm for iteratively optimizing the prompt by finding failing examples and adding to the prompt.
-But would such a procedure converge to a good solution?
-We can view $f(\mathbf{x})$ as a set function over the examples (that compose $\mathbf{x}$).
-A natural property to examine is submodularity (where greedy optimization provably converges to a good solution under some conditions [15]), i.e., the property of diminishing marginal utility of examples.
-The marginal utility of adding an example $e$ to the few-shot prompt corresponding to a set of examples $A$ is $\Delta_{A}=f(A\cup\{e\})-f(A)$.
-Let $B\subset A$ denote a subset of $A$.
-We define $f_{\text{fewshot}}$ to be the validation accuracy of GPT-3.5 using few-shot prompts on the Ethos dataset (see Section 4).
-In Figure 2 (b), we show the distribution of diminishing marginal utilities for $f_{\text{fewshot}}$ (light blue), i.e., we sample multiple $A,B$ sets of examples from the dataset and compute $\Delta_{B}-\Delta_{A}$, which is expected to be non-negative for submodular functions.
-As we see, the marginal utility differences are nearly uniformly random in $[-0.01,0.01]$, indicating that there is no submodularity structure to $f_{\text{fewshot}}$.
-A recurring theme [1, 2] in the prompts that humans engineer is to organize task information in “sections” (introduction, context, etc.), treating prompt like a structured document.
-When humans engineer a prompt, they do it systematically by editing specific sections of the prompt.
-We hypothesize that there are a small number of facets that collectively define a task, which can be discovered by inspecting a representative set of examples from the data distribution.
-Consider the task of tagging hate speech.
-Identifying intent to harm, recognizing general aggression versus targeted hate, and differentiating hate from other objectionable content are some facets of the task.
-We ask whether there is submodularity structure in adding different facets (text sections) to a prompt, compared to adding in-context examples.
-The answer to this question depends on multiple factors: (1) what are the contents of the sections and how are they generated?; (2) how does $f(\mathbf{x})$ behave w.r.t.
-the sections that make up the prompt $\mathbf{x}$?
-Figure 2 (b) shows the distribution of marginal utility differences for $f_{\text{facets}}$ (dark blue), i.e., we sample multiple $A,B$ sets of facets (i.e., short sections) that are generated using a Llama2-13B model, and compute $\Delta_{B}-\Delta_{A}$ (defined as above, but the added element $e$ is now a section) using GPT-3.5 as the solver LLM (See Appendix A.2 for experimental details).
-We see that the diminishing marginal returns property largely holds for $f_{\text{facets}}$, unlike in the case of $f_{\text{fewshot}}$.
-This suggests that greedy strategies for prompt optimization problem over sections could work — but this crucially relies on how we generate the sections themselves.
-The prompt optimization problem can now be posed as one of inferring the facets of a task from a set of training examples.
-In the following section, we propose a method that iteratively learns the facets of the task as sections of the prompt.
-3 UniPrompt: Generating a Prompt that Captures Multiple Facets of a Task
-Manual prompt engineering consists broadly of two steps: 1) generating an initial prompt; and 2) identifying errors and editing the prompt.
-The initial prompt for a task is constructed based on domain knowledge.
-Typically, the initial prompt contains implicit or explicit sections (e.g., introduction, task description, intended application, edge-cases, challenges, etc.) wherein each section corresponds to a relevant task facet.
-Given the initial prompt and a solver LLM, a human prompt engineer would look at the incorrect predictions, form hypotheses on which section of the prompt needs to be edited or whether a new section needs to be added, and then propose edits that may fix the errors.
-Then, the prompt engineer may pick the edit with the highest validation accuracy.
-These steps are repeated in a loop, until the validation accuracy saturates.
-A good engineer may also backtrack from an optimization trace and start from some previous prompt, in case the current optimization trajectory yields unsatisfactory results.
-We present UniPrompt, a prompt optimization algorithm that aims to mimic the above process.
-Specifically, the algorithm extracts key concepts or facets relevant to a task and updates prompt sections using them, with the overall goal of increasing validation accuracy.
-Our analysis from Section 2 indicates that such a directional optimization procedure based on iteratively adding task facets can converge to a good solution, especially for solver LLMs such as GPT-3.5 and GPT-4.
-[…]
-1) Task Facets Initialization using background knowledge.
-In the first stage, the algorithm leverages a language model to generate a prompt with multiple facets, i.e., sections focusing on specific task aspects based on the task description.
-It greedily selects content for each facet from a pool of candidates generated using the language model.
-2) Task Facets Refinement using examples.
-In the second stage, the algorithm refines the prompt using a set of training examples and an auxiliary LLM.
-Given a candidate prompt, it collects examples with prediction errors by the solver LLM, extracts key concepts or facets required to fix those errors, and makes appropriate edits.
-An edit may create a new section, modify an existing section, or delete an existing section.
-Additionally, we maintain a beam of best performing prompts to backtrack and restart from a previous prompt if the current optimization path is unsatisfactory.
-Similar to UniPrompt, previous methods [12, 22] propose using an auxiliary LLM to provide feedback on incorrect predictions and propose edits to a given prompt.
-In our initial experiments, however, we find that collecting feedback over individual examples or randomly sampled batches leads to memorization of individual examples (see Figure 1) rather than recognition of facets that are important to the task.
-To avoid overfitting to the train examples and to extract generalizable concepts, we make two contributions.
+1.
+Larger models are more amenable to prompt optimization.
+We observe that the change in the objective function (i.e., loss on a validation set for a given prompt) per change in input is relatively more stable for larger models like GPT-4 (Figure 2) than for GPT-3.5 (analysis in Appendix A.1).
+2. Clustered-batching improves the quality of text gradients (i.e., feedback), as against the standard random batching adopted in state-of-the-art prompt optimization methods (Section 5.1).
+3.
+Two-tier feedback helps learn generalizable facets.
+Collecting feedback from an expert LLM over mini-batches, and then summarizing the individual feedback texts via a second step (Section 5.1) helps learn generalizable task concepts in prompts.
+The proposed method UniPrompt, in Figure 1 [Right], makes two contributions.
 First, we follow a two-tier setup of synthesizing feedback for a batch of training examples.
 We break up a batch into mini-batches, collect feedback on each of the mini-batches and then use a separate prompt to aggregate the different feedback texts into a generalizable concept.
-Second, to increase chances that a mini-batch corresponds to a coherent facet, we cluster the training data beforehand and ensure that each mini-batch consists of examples from the same cluster.
-We provide details of the algorithm below (see Algorithm 1).
-Notation: We denote the training set with $D_{t}$ where each example is a question-answer pair of the form $\{q_{i},a_{i}\}$.
-The validation set is denoted as $D_{v}$.
-We have $N$ training examples and $K$ validation examples.
-Input to the algorithm is the solver LLM $f_{LLM}$, train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$.
-In addition, we assume access to an “expert” auxiliary LLM such as GPT-4.
-Input: Train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$
-Output: Optimized prompt $P^{*}$ for the given task
-1 Cluster train set, initialize history and validation accuracy arrays: $C\leftarrow$ cluster($D_{t}$), $H\leftarrow\{\}$, $V\leftarrow$ [];
-2 Initialize a beam of size 2 with the initial prompt: $p_{1}\leftarrow p_{0}$ and $p_{2}\leftarrow p_{0}$ for each cluster c in C do
-3       $B\leftarrow$ batches($C$);
-4       for each batch $b\in B$ do
-5             $M\leftarrow$ mini-batches($B$)
-6             $F\leftarrow[]$
-7             for each mini-batch $m\in M$ do
-8                   Evaluate the best prompt on mini-batch: $a_{m}\leftarrow$ LLM($m,p_{1}$)
-9                   Get feedback from the expert given history of mini-batch, accuracy and task description: $f\leftarrow$ Feedback($T$, $a_{m}$, $H[m]$)
-10                   $F$.insert($f$)
-11
-12            Combine feedbacks over a batch: $F_{b}\leftarrow$ Combine($F$)
-13             Apply feedback to get updated prompts; $q_{1}\leftarrow$ apply($F_{b}$, $p_{1}$); $q_{2}\leftarrow$ apply($F_{b}$, $p_{2}$)
-14             Update the beam: if not($p_{1}$ = $p_{0}$) then
-15                   $p_{2}\leftarrow$ second-high-acc([$p_{1}$, $p_{2}$, $q_{1}$, $q_{2}$], b)
-16            $p_{1}\leftarrow$ highest-acc([$p_{1}$, $q_{1}$, $q_{2}$], b)
-17      Evaluate the best prompt on validation set: $acc_{v}\leftarrow$ evaluate($p_{1}$, $D_{v}$)
-18       $V\leftarrow V$.append($acc_{v}$)
-19       Stop based on the validation accuracy: $c\leftarrow$ early-stop-criteria($V$)
-20       if $c$ then
-21             break
-22
-23return $p_{1}$;
-Algorithm 1 UniPrompt
-3.1 Task facet initialization using background knowledge
-While existing methods for prompt optimization [12, 22] iteratively edit a human-written prompt, our goal is to generate the entire prompt from scratch (specifically, based only on a one line description of the task).
-Therefore, one option is to initialize the prompt using only the task description.
-We call this Task Description Initialization and initialize the prompt with a single section titled Introduction containing the input task description.
-In the second kind of initialization, we finetune a language model to generate a prompt with multiple diverse sections such as Introduction, Tricks, and Corner Cases; similar to the initial prompt that a human prompt engineer may produce.
-Here we use the world knowledge embedded in language models that are trained on extensive text data.
-The language model is finetuned to output the entire section contents given the task description and the section title.
-Specifically, we train a Llama2-13B model using a synthetic dataset generated using GPT-4.
-Across a wide range of tasks, we first use use GPT-4 to identify the important sections given the task, section descriptions and a few examples (See Appendix A.6 for the prompt).
-We then fine-tune a language model $\phi$, that is Llama2-13B, in a supervised setting on a synthetically created dataset of the form $\langle T,P_{i},S_{i},{\bf S^{\prime}}_{\text{gold}}\rangle$ where $T$ denotes the task description, $P_{i}$ is the prompt generated using the fine-tuned model sequentially before section $S_{i}$ (i.e., it would have optimized sections $\{S_{1},S_{2},...S_{i-1}\}$), and ${\bf S^{\prime}}_{\text{gold}}$ is the gold section contents.
-$\phi$ is finetuned on the objective: $\min_{\phi}[-\log(p_{\phi}({\bf S^{\prime}}_{\text{gold}}|T,P_{i},D_{t},S_{i}))]$.
-We then sample 10 candidates for each section using the fine-tuned model and evaluate each section added incrementally against $D_{v}$ and select the best performing candidate for each section.
-More details on data collection and training of SLM are given in the Appendix A.4.
-Examples of one line task description and the Llama-initialized prompt for different tasks are in Appendix A.3.
-3.2 Task facet refinement using examples
+Second, to increase chances that a mini-batch corresponds to a coherent facet, we periodically (re)cluster the training data and ensure that each mini-batch consists of examples from the same cluster.
+Algorithm 1 receives as input a one-line task description and a train set $D_{t}$ of $N$langlequestion $q_{i}$, answer $a_{i}\rangle$ demonstrations.
+It extracts key concepts or facets relevant to the task and updates prompt sections using them, with the goal of increasing accuracy on the validation set $D_{v}$.
+We assume access to an “expert LLM” such as GPT-4.
+3.1 Task facet learning using examples
 Extracting task-relevant concepts from a set of examples to refine a prompt is a complex problem comprising multiple steps.
-Given a set of incorrect predictions, one needs to analyze what went wrong in each prediction, form hypotheses, aggregate the hypotheses to identify specific concepts that are relevant for the task, and then for each concept, one needs to attribute which facet/section of the current prompt needs to be edited to incorporate the concept.
+Given a set of incorrect predictions, one needs to analyze what went wrong in each prediction, form hypotheses, aggregate the hypotheses to identify specific concepts that are relevant for the task.
+Then, for each concept, one needs to attribute which facet/section of the current prompt needs to be edited to incorporate the concept.
 These operations are highly model-specific and are difficult to execute reliably.
-Therefore, for this stage, we exclusively rely on an “expert” auxiliary LLM such as GPT-4.
-The whole process is executed through a series of prompts to the expert LLM.
+Therefore, we exclusively rely on an expert LLM.
 First, we prompt the expert LLM to diagnose mistakes (feedback) in each example given the answer and chain-of-thought reasoning produced by the solver LLM.
@@
 To aid in identifying major edits that correspond to generalizable facets, we propose to cluster the examples as a preprocessing step and create clustered batches, such that each cluster shares some common facet of the task.
-3.2.1 Preprocessing step: Clustering examples to facilitate facet identification
-We explore two approaches for clustering: topic based clustering, and feedback based clustering.
+3.1.1 Clustering for identifying facets
+We explore two approaches for clustering: topic-based clustering, and feedback-based clustering.
 Topic-Based Clustering.
-Given a set of examples, we identify $l$ topics that would span the entire train set.
+Given a set of examples, we identify $l$ topics spanning the entire train set.
 This type of clustering is motivated by the observation that solver LLM may make similar mistakes on examples from the same topic.
@@
 Feedback-Based Clustering.
-Another way to find examples that share similar task facets is the feedback they receive based on the initial prompt’s predictions.
-For instance, for a physics-based task, if two examples from different topics obtain the same feedback from the expert LLM to edit the “Rules” section of the prompt and include the statement, “Draw all forces on each body before writing the equations", then we argue that such examples can be clustered.
+Examples that receive similar feedback based on a prompt’s predictions can help identify task facets.
+Consider a physics-based task where two examples from different topics obtain the same feedback from the expert LLM to edit the “Rules” section of the prompt to include the statement, “Draw all forces on each body before writing the equations”.
+We argue that such examples can be clustered.
 This type of clustering makes the broad edit identification step easier.
-To obtain the clusters, we first evaluate all training examples against the initial prompt $p_{0}$ and store the feedback $f_{i}$ from the expert LLM, corresponding to each incorrectly answered example ${q_{i},a_{i}}$ (all the correctly answered questions form one cluster).
-We then prompt the expert LLM to cluster these feedbacks $\{f_{1},f_{2},\dots,f_{N}\}$ into $l$ clusters $\{f^{\prime}_{1},f^{\prime}_{2},\dots,f^{\prime}_{l}\}$ (see Appendix A.8).
-For each cluster $f^{\prime}_{j}$, we create a batch of examples ${q_{i},a_{i}}$ corresponding to each $f_{i}$ in $f^{\prime}_{j}$.
-3.2.2 Generating prompt edits that generalize to multiple examples
+To obtain the clusters, we first evaluate all training examples against the current best prompt and store the feedback $f_{i}$ from the expert LLM, corresponding to each incorrectly answered example ${q_{i},a_{i}}$ (all the correctly answered questions form one cluster).
+We then prompt the expert LLM to cluster these feedbacks $\{f_{1},f_{2},\dots,f_{N}\}$ into $l$ clusters (see Appendix A.11).
+For each cluster, we create a batch ${q_{i},a_{i}}$ corresponding to feedbacks in that cluster.
+3.1.2 Obtaining generalizable prompt edits
 Two-tier Feedback.
-To encourage generalizable feedback from the expert LLM that applies to more than one example, we obtain feedback at two levels: mini-batch and batch.
+To encourage generalizable feedback from the expert LLM, we obtain feedback at two levels: mini-batch and batch.
 Given a batch (created using clustering discussed above), we break it up into mini-batches.
-For each mini-batch $m$, we consider the questions for which the solver LM answered incorrectly with the current prompt.
-For these questions, we construct a prompt consisting of the questions, the chain-of-thought produced by the solver LLM, their incorrect predicted answers and the ground-truth answers.
+For each mini-batch $m$, we construct a prompt consisting of incorrectly-answered questions in $m$, the chain-of-thought produced by the solver LLM, their incorrect predictions and the ground-truth answers.
 We ask the expert to provide one feedback for the mini batch (prompt is provided in Appendix A.12).
-In the prompt, we specify the following types of edits that the expert LLM can suggest: add a section or subsection, delete a section or subsection, and edit a section or subsection.
-Given the different feedback texts for mini-batches within a batch $b$, the next step is use another call to the expert LLM to summarize the given set of edits and provide a single summarized section update.
+The expert can suggest the following edits: add a section or subsection, delete a section or subsection, and edit a section or subsection.
+Given the different edits for mini-batches within a batch $b$, we invoke the expert LLM again to summarize these edits into a single section update.
 This combination ensures some degree of smoothness at every update which helps stabilize training.
@@
 As before, the class of edits allowed is the same.
-History for effective exploration: To ensure comprehensive and non-repetitive exploration of prompts, we also provide the batch-level history of edits in the mini-batch-level prompt.
-Specifically, we provide the expert with the history of edits $H[b]$ [12, 31] for the batch $b$ and the corresponding accuracy difference that each edit caused over the previous edit for the given batch.
-History is presented in the format $\{e_{i},acc_{i}-acc_{i-1}\}$ where $e_{i}$ is the edit proposed at the $i^{th}$ update and $acc_{i}$ is the accuracy of the $i^{th}$ updated prompt (See Appendix A.9 for the full prompt).
+History for effective exploration.
+To ensure comprehensive, non-repetitive exploration of prompts, we also provide the batch-level history of edits (Hu et al., 2024; Yang et al., 2023) in the mini-batch-level prompt.
+History $H[b]$ is presented as $\{e_{i},acc_{i}-acc_{i-1}\}$ where $e_{i}$ is the edit proposed at the $i^{th}$ update and $acc_{i}$ is the accuracy of the $i^{th}$ updated prompt (See Appendix A.12 for the prompt).
+Input: Train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$
+Output: Optimized prompt $P^{*}$ for the given task
+1 $C\leftarrow$ cluster($D_{t},p_{0}$), initialize history $H\leftarrow\{\}$, and validation accuracies $V\leftarrow$ [];
+2 Initialize a beam of size 2 with the initial prompt: $p_{1}\leftarrow p_{0}$ and $p_{2}\leftarrow p_{0}$
+3 for epoch $e$ and each cluster c in C do
+4
+5      for each batch $b\in\texttt{batches}($C)))) do
+6             $F\leftarrow[]$
+[…]
+8                   Evaluate the best prompt on mini-batch: $a_{m}\leftarrow$ LLM($m,p_{1}$)
+9                   Get expert feedback: $f\leftarrow$ Feedback($T$, $a_{m}$, $H[m]$)
+10                   $F$.insert($f$)
+11
+12            Combine feedbacks over a batch: $F_{b}\leftarrow$ Combine($F$)
+13             Apply feedback to get updated prompts: $q_{1}\leftarrow$ apply($F_{b}$, $p_{1}$); $q_{2}\leftarrow$ apply($F_{b}$, $p_{2}$)
+14             Update the beam: if not($p_{1}$ = $p_{0}$) then $p_{2}\leftarrow$ second-high-acc([$p_{1}$, $p_{2}$, $q_{1}$, $q_{2}$], b)
+15             $p_{1}\leftarrow$ highest-acc([$p_{1}$, $q_{1}$, $q_{2}$], b)
+16      Evaluate the best prompt on validation set: $acc_{v}\leftarrow$ evaluate($p_{1}$, $D_{v}$)
+17       $V\leftarrow V$.append($acc_{v}$)
+18       if early-stop-criteria ($V$) then break
+19       if recluster($e$) then $C\leftarrow$ cluster($D_{t},p_{1}$)
+20return $p_{1}$ as $P^{*}$;
+Algorithm 1 UniPrompt
 3.1.3 Editing the prompt
-Once the final set is edits is received for a batch, we use the expert LLM to apply edits to the current prompt (see Appendix A.10 for the prompt).
-An edit is accepted only if it increases the validation accuracy compared to the current prompt.
-We call this approach “Greedy”.
-Alternatively, we maintain a beam of $l$ best performing prompts based on validation accuracy, apply the edit to all the $l$ prompts, and update the beam to retain the top $l$ performing prompts.
-We call this method “Beam”.
-To avoid overfitting on the train examples (or keep adding unnecessary information to the prompt), we employ early stopping in the optimization process (More details given in Section 4).
-3.3 Efficiency considerations: Computational Complexity
-The computational complexity of UniPrompt is dominated by the training phase, and in particular, by the queries made to the solver and the expert LLMs; at inference time, we need a single query to the solver LLM using the learnt prompt and the given test example.
-This is particularly well-suited for settings such as recommendation systems, where inference time efficiency is critical.
-In contrast, methods such as MedPrompt [21] require multiple calls at test time for a single test example.
-Let us now consider the compute complexity in terms of the number of expert or solver LLM calls made per epoch, stage-wise.
-Clustering: First, we evaluate all the training examples using the current prompt.
-Second, for every wrongly predicted example, we obtain feedback from the expert LLM.
-Third, for the given set of feedbacks, we use a single call to cluster it into $l$ clusters.
-Each of the above steps incurs $O(N)$ queries, so the total query complexity of the clustering stage is $O(N)$.
-Finally, for each example, i.e., (question, answer) pair, we simply map it to the $l$ clusters (no LLM calls).
-This is a one-time cost.
-Mini-batch feedback and Batch-level aggregation: At a given epoch, we evaluate every question in the mini-batch using the current prompt and the solver LLM ($N$ queries overall).
-Next, we obtain one feedback over all the wrong questions in the mini-batch $m$ ($N/|m|$ queries).
-We use one call to aggregate these feedbacks.
-For prompt selection, we evaluate $2k$ prompts on the batch $b$, where $k$ is the beam size ($O(2k|b|)$ queries per batch).
-Hence overall query complexity is $N+N/|m|+2kN+1$ or $O(kN)$.
-In practice, $k$ is a small number (2 in our experiments).
-Assuming LLM throughput of 0.5 queries per second, a training + validation set of 300 examples, $l=10$ clusters, beam size $k=2$ and 20 epochs, it takes under 7 hours to train.
-4 Experiments
-Datasets We perform evaluation on four standard datasets : (1) Ethos [20], (2) ARC [4], (3) MedQA [13], and (4) GSM8K [5].
+Once the final set of edits is received for a batch, we use the expert LLM to apply edits to the current prompt (See Appendix A.13 for the prompt).
+An edit is accepted only if it increases the validation accuracy (Greedy).
+Alternatively, we maintain a beam of 2 best performing prompts based on validation accuracy, apply the edit to the two prompts, and update the beam to retain the top 2 performing prompts (Beam).
+To avoid overfitting on the train examples (or adding unnecessary information to the prompt), we employ early stopping in the optimization process (more details in Section 4).
+3.2 Prompt Initialization
+We use two types of initialization: (1) task description, i.e., $p_{0}$ has a single section titled Introduction containing the input task description.
+(2) finetune Llama2-13B model to generate a prompt with sections such as Introduction, Tricks, and Corner Cases, similar to the initial prompt that a human prompt engineer may produce.
+To finetune, we use GPT-4 generated data consisting of (task description, section title, section contents) triples.
+Details and examples are in Appendices A.4 and A.7.
+Computational Complexity:
+The complexity of clustering and of getting mini-batch and batch-level feedbacks per epoch is $O(N)$ expert LLM queries, where $N$ is the number of training examples. Details are in Appendix A.3.
+4 Experiments Setup
+Datasets: We perform comprehensive evaluation on five standard datasets: (1) Ethos (Mollas et al., 2020), (2) ARC (Clark et al., 2018), (3) MedQA (Jin et al., 2021), (4) GSM8K (Cobbe et al., 2021), and (5) BBH (Suzgun et al., 2022).
 Ethos, ARC, and MedQA contain multiple choice questions, and GSM8K contains questions with integer answers.
-Implementation details We set the initial prompt $p_{0}$ for each task as the one line task description and use 200 data points $(q_{i},a_{i})$ as the train set.
-We use GPT-3.5-Turbo as the solver model, that is used in a chain-of-thought setting.
-For Feedback and Combine, we use GPT-4 as the expert.
+BBH is a subset of 10 tasks, spanning 4 main categories, from the challenging BIG-Bench benchmark that requires multi-step reasoning.
+In addition, we also evaluate UniPrompt on the medical QnA datasets used in the MedPrompt (Nori et al., 2023) work; as well as two popular code generation datasets, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
+Implementation details: We set the initial prompt $p_{0}$ for each task as the one-line task description.
+We use 200 examples as the train set, 200 examples as the test set, and 100 examples as the validation set for all the compared methods.
+We use GPT-3.5-Turbo as the solver model.
+For Feedback and Combine in UniPrompt, we use GPT-4 as the expert (see ablation in Section 5.5).
 We maintain a beam size of 2.
@@
 We find that mini-batch sizes 3 to 5 and batch sizes 5 to 7 work the best for our datasets.
-The temperature of the LLMs for our method is set to 0 for reproducibility of results.
+The temperature of the LLMs for our method is set to 0 for reproducibility.
 We employ early stopping at batch-level in UniPrompt.
-We find that a $patience$ value of 5 works best for the given tasks.
-Baselines We compare UniPrompt with the following techniques: (1) Task Description: prompt is the one line task description that we use to initialize UniPrompt; (2) Chain-Of-Thought (or CoT) prompting [14]; (3) Expert Prompt: the prompt optimized by humans taken from prior works [21]; (4) OPRO [31], that uses LLMs for discrete optimization over text prompts; (5) ProTeGi [22] that proposes textual gradients and selects edits to prompts using bandit techniques; (6) Evoke [12] that uses two instances of LLM, one that scores the current prompt, and the other that edits the prompt; (7) EvoPrompt [7] that uses genetic algorithms to search through the space of prompts; and (8) MedPrompt [21] that effectively combines multiple prompting strategies such as kNN-based few-shot selection, CoT reasoning, and ensembling.
-Metrics For each dataset, we perform 5-fold validation (we split the train set into 5 folds) and report the final test accuracy as the average accuracy of the 5 runs on the test sets.
-4.1 How does UniPrompt compare to existing methods in zero-shot setting
-In the zero-shot setting, we do not include labeled examples in the prompt for any of the compared methods.
-We report results for four versions of our method in Table 1; they differ in the initial prompt—task description, expert prompt, or the prompt generated by the fine-tuned Llama model (discussed in Section 3.1)—and the combining strategy (from Section 3.2.3)—beam search vs greedy.
+Baselines: We compare UniPrompt with the following techniques and baselines: (1) Task Description: prompt is the one line task description that we use to initialize UniPrompt; (2) Chain-Of-Thought (or CoT) prompting (Kojima et al., 2024); (3) Expert Prompt: the prompt optimized by humans taken from prior works (Nori et al., 2023); (4) OPRO (Yang et al., 2023), that uses LLMs for discrete optimization over text prompts; (5) ProTeGi (Pryzant et al., 2023) that proposes textual gradients and selects edits to prompts using bandit techniques; (6) Evoke (Hu et al., 2024) that uses two instances of LLM, one that scores the current prompt, and the other that edits the prompt; (7) EvoPrompt (Guo et al., 2023) that uses genetic algorithms to search through the space of prompts; (8) TextGrad (Hou et al., 2023), state-of-the-art framework for automatic differentiation of prompts via text; (9) DSPy (Khattab et al., 2024), a recent programming model for optimizing LLM prompts; and (10) MedPrompt (Nori et al., 2023), a state-of-the-art prompt composition method.
+5 Results and Analysis
+We present detailed quantitative and qualitative results, along with key ablations.
+5.1 Performance of UniPrompt
+We start with the zero-shot setting, where we do not include labeled examples in the prompt for any of the compared methods.
+We report results for two versions of our method in Table 1, which differ in the combining strategy (from Section 3.1.3)—beam search vs greedy.
 UniPrompt variants significantly outperform the baselines including CoT and the state-of-the-art prompt optimization techniques like ProTeGi that crucially leverage LLMs for performing iterative prompt edits.
-We achieve maximum gains on the Ethos dataset with a $18.2\%$ increase in accuracy over the expert prompt.
-Further, we see accuracy increases of $4.0\%$ on MedQA, $3.5\%$ on GSM8k, and $7.6\%$ on ARC-Challange datasets.
-An example of evolution of prompts using our algorithm is given in Appendix4.
-It starts with a simple description of task and adds important details like differentiating between hate speech and rudeness.
-In contrast, ProTeGi [22] yields a rather terse prompt on the same dataset: “Does the following text contain language that targets a group of people based on their religion, gender, or other personal characteristics?”.
-This prompt does not provide various facets of the task, the way our method infers or an expert would typically write.
-Table 1: Test accuracies for the compared methods with GPT-3.5-Turbo as the solver model in the zero-shot setting (best in bold; second best underlined).
-The last four rows correspond to our proposed method.
+UniPrompt is the best performing method on three out of four datasets.
+It achieves maximum gains on the Ethos dataset with a $18.2\%$ increase in accuracy over the expert prompt.
+Further, we see accuracy increases of $4.0\%$ on MedQA, $3.5\%$ on GSM8k, and $7.6\%$ on ARC-Challenge datasets.
+We show UniPrompt training behavior in Appendix A.19.
+We also present comparisons to state-of-the-art DSPy method in the few-shot setting (8 bootstrapped_demos) using two optimization settings provided by their framework.
+The last two rows of Table 1 show that UniPrompt in the zero-shot setting convincingly outperforms DSPy in the few-shot setting, on three out of four datasets.
+Table 1: Test accuracies (%) of the compared methods with GPT-3.5-Turbo as the solver model in the zero-shot setting (best in bold; second best underlined).
+The two UniPrompt rows are our proposed method.
+We compare with few-shot methods in the last two rows (DSPy variants); *best in bold to distinguish the few-shot setting.
 Method Ethos ARC MedQA GSM8K
@@
 EvoPrompt 81.6 89.9 50.3 81.4
-UniPrompt (Init = Expert Prompt) + Beam 84.0 86.0 52.3 82.4
+DSPy (MIPRO v2, zero-shot) 79.7 82.8 61.9 77.3
+TextGrad 79.5 76.5 50.6 81.6
 UniPrompt (Init = Task Description) + Beam 92.3 86.0 57.1 82.4
-UniPrompt (Init = Llama Prompt) + Beam 92.0 90.5 55.5 81.5
 UniPrompt (Init = Task Description) + Greedy 93.7 90.5 55.5 82.3
-The training curves in Figure 3 show that our method initially performs edits on the prompt that simultaneously increase the train as well as the validation accuracy.
-After about 10 or 15 iterations (each batch update is an iteration), validation accuracy decreases while train accuracy continues increasing, indicating overfitting; which we overcome by using early stopping.
-4.2 How does UniPrompt perform compared to SOTA in the few-shot setting?
-We now compare the performance of UniPrompt, in the few-shot setting with GPT-4 as the solver model, to the SOTA prompting technique MedPrompt [21] that employs three key components: (1) few-shot prompting, where they dynamically select five relevant examples using k-nearest neighbors (kNN) clustering in the embedding space; (2) Chain-of-Thought (CoT) reasoning on the selected examples; and (3) self-consistency and ensembling with option shuffling.
-From Table 2, we observe that UniPrompt (first col), which requires only one call at inference time, performs almost as well as MedPrompt (last col), which requires five calls, on the MedQA, PubMedQA, and MedMCQA datasets.
-As we incrementally add components such as kNN few-shot, CoT, and Ensemble to our prompt, we see a significant increase in accuracy of $4.35\%$ on average across all datasets.
-4.3 Results on a Real-world Task: Search Query Intent
-We consider the task of inferring if two search queries share identical intent or not.
-This is an important task arising in search and recommendation pipelines, and is tackled today using LLMs.
-It is challenging to tell if two queries share the same intent, as it requires domain knowledge (e.g., brands and product categories), depends on geographical biases (e.g., “cricket” and “cricket game” are likely to mean the same in UK, but unlikely in the US), and even on cultural connotation.
-So, examples are crucial for understanding the task and engineering a prompt that generalizes well.
+DSPy (BootstrapFewShotWithRandomSearch) 86.6 87.5 *68.5 74.3
+DSPy (MIPRO v2, few-shot) 84.0 86.0 62.9 79.7
+Qualitative Analysis: ProTeGi and TextGrad also adopt batching by randomly sampling from training examples where the solver LLM made mistakes.
+In the early iterations of optimization, there can be many such examples.
+So, do our key observations and hypotheses (beginning of Section 3) hold empirically?
+We give some evidence below.
+1. Employing clustering to create batches (Section 3.1): An example feedback obtained on the Ethos dataset using UniPrompt is shown below:
+The instruction should include..potential harm or violence implied, as well as any discriminatory or derogatory language used…towards a particular religious group.
+The instruction should include …think about the impact of the statement on the targeted individual or group.
+The instruction should…language that prompts hatred or discrimination towards a particular gender.
+To contrast, we employ random batching as in the standard prompt optimization techniques, on the same dataset.
+The feedback obtained, given below, is relevant for the task, but fails to identify specific concepts.
+The instruction should include a clear definition of hate speech…
+The instruction should include examples of hate speech, guidance on identifying hate speech…
+The former feedback (UniPrompt) captures the facet of measuring impact on the targeted entity whereas the latter only captures religious and harm-based aspect of hate speech.
+The same expert LLM is able to identify different facets due to clustered batches.
+2.
+Employing two-tier feedback: In Section 3, we argued that employing two-tier feedback strategy to aggregate the feedback texts encourages the expert LLM to propose edits that are generalizable.
+The following feedback is received on the Ethos dataset after the aggregation:
+The instruction should…consider whether the statement contains discriminatory, derogatory, or violent language that promotes hatred or harm towards a particular group, such as based on religion and gender.
+We see that two-tier feedback helps in distilling important aspects of the task implicit in the examples, rather than directly using or rephrasing the (limited) examples.
+Results on BBH: From Table 2, it is evident that UniPrompt shows a significant improvement over OPRO (that also evaluates on these tasks in their paper) for a majority of tasks.
+It achieves significantly higher accuracy in Boolean Expression (92.37% vs.
+78.74%), Date Understanding (81.96% vs.
+52.59%), and Navigate (77.16% vs.
+51.74%).
+Table 2: Test accuracies (%) on BBH dataset with GPT-3.5-Turbo as the solver model
+Task Init OPRO UniPrompt
+Algo & Multi-Step Arithmetic Reasoning
+Bool Exp. 83.64 78.74 92.37
+Logical Ded. 29.53 38.97 39.62
+Navigate 60.95 51.74 77.16
+Natural Language Understanding
+Snarks 67.00 67.88 74.30
+Disamb. QA 53.30 57.43 67.05
+Fallacies 57.60 53.14 57.90
+Use of World Knowledge
+Causal Judg. 54.29 57.24 59.37
+Movie Rec. 58.04 77.81 71.80
+Dates 74.21 52.59 81.96
+Multilingual Knowledge & Reasoning
+Salient Trans. 42.59 50.61 50.77
+5.2 Comparison with MedPrompt
+MedPrompt (Nori et al., 2023) is a recent, competitive prompting technique without any training component.
+It employs three key ingredients: (1) few-shot prompting, where five relevant examples are selected using k-nearest neighbors (kNN); (2) CoT reasoning on the selected examples; and (3) self-consistency and ensembling with option shuffling at inference time.
+They evaluate on 4 medical datasets (that none of the competing methods in Table 1 evaluate on) using GPT-4 as the solver model.
+So, we compare UniPrompt in the same setting in Table 8 (in Appendix A.8).
+UniPrompt (first row), which requires only one call at inference time, performs almost as well as MedPrompt (last row), which requires five calls, on three out of four datasets.
+As we incrementally add kNN few-shot, CoT, and ensembling to our prompt, we see a significant increase in accuracy of $4.35\%$ on average across all datasets.
+5.3 Performance on generation tasks
+Our evaluations so far have been on multiple-choice QnA, math, and classification datasets.
+We now evaluate UniPrompt on generating code given a natural language specification.
+We use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) datasets consisting of Python coding problems.
+We initialize with a simple prompt, “You are a software engineer.
+You are given a function signature and a description of the function.
+You have to complete the function.” We use GPT-4-Turbo as both the solver and the expert LLM.
+HumanEval does not have train or validation sets.
+So, we take random 100 examples from MBPP as train.
+Similarly, for MBPP, we take random 50 examples from HumanEval as train.
+We evaluate the final prompts on HumanEval and MBPP test sets.
+The results are given in Table 4 (in Appendix A.8).
+The metric is % of solved coding problems (evaluated using the provided test cases) in the datasets.
+The prompts produced by UniPrompt outperform standard prompting of LLMs.
+5.4 Results on a Real-world Task
+The task of inferring if two search queries share identical intent or not arises in search and recommendation pipelines.
+It is challenging because it requires domain knowledge (e.g., brands and product categories), and depends on cultural and geographical biases (e.g., “cricket” in UK vs.
+“cricket game” in the US).
+So, examples are crucial for engineering a prompt that generalizes well.
+We sample 200 train, 50 validation, and a separate 2527 test user queries from a real proprietary application. More details on the dataset and the one-line initial prompt are provided in Appendix A.5.
+The prompt obtained using UniPrompt improves over the best manual prompt by 5.77% on the negative (rare) class, by 0.23% on the positive class, and by 1.86% overall on the test set.
+The learnt prompt captures the following task facets: (1) recognizing variations in names and abbreviations, and how they do not change the context; (2) recognizing brand specificity, and how even minor variations do change the context; and (3) recognizing the specificity of terms in queries, and how lack of specific terms can indicate departure of intent.
+5.5 Ablations
+Impact of Clustering, Inclusion of History, and Greedy Update: The results are shown in Table 6 (in Appendix A.8).
+We see that clustering as well as edit history components (Section 3.1) are critical for performance of UniPrompt in all the datasets.
+We see a major drop of $14.8\%$ in accuracy in the Ethos dataset when clustering is removed, and a $4.3\%$ drop when history component is removed.
+In all the datasets except GSM8K, we find clustering is more important than history.
+This can attributed to limited variability of question types (all grade-8 arithmetic) in GSM8K than in others.
+We also find that the greedy update rule (Section 3.1.3) proves to be superior or competitive compared to beam search in relatively easier datasets — where even less exploration produces good results, greedy proves to be a more effective update rule.
+On the other hand, in more complex datasets like MedQA, greedy appears to be a bad strategy.
+We also see that clustering examples based on feedback (“Fb Clustering”) is a better strategy than clustering based on topics, except for the Ethos dataset.
+Impact of Mini-batch size and Number of Clusters:
+Table 3: Impact of UniPrompt’s key hyperparameters
+Hyperparameter Value Accuracy
+Mini-batch Size 2 84.90%
+5 90.36%
+8 91.15%
+Number of Clusters 2 85.81%
+5 90.36%
+10 87.82%
+First, we vary mini-batch size in UniPrompt.
+The results for the ARC dataset are shown in Table 3 (in Appendix A.8).
+With an increase in mini-batch size, we observe an increase in accuracy.
+That said, it is a hyperparameter, hence there will be an optimal number for each dataset.
+The mini-batch size affects the feedback based on the wrong examples that are obtained in each round.
+Next, we vary the number of clusters in UniPrompt.
+We find that the parameter has a clear impact on performance.
+We use the default choice of 5 clusters in all our experiments, which provides concise and generalizable feedback.
+Impact of initial prompt and Expert model capacity: In Table 7 (Appendix A.8), we find that one-line task description initialization for UniPrompt achieves the best accuracy on three out of four datasets.
+On ARC, initializing with the prompt generated by the Llama2-13B model gives significant improvement over other initializations.
+In Table 5 (Appendix A.8), we show UniPrompt improves prompts for more capable solver LLMs while using less capable expert LLMs.
+6 Conclusions
+We presented a method inspired by the human prompt engineering process to generate complex prompts from scratch that include different facets of a task.
+Our algorithm provides significant improvements over baseline prompt generation methods on multiple standard datasets.
+Just like in-context learning (Ji et al., 2024), task facet learning could also benefit from connections to submodular optimization (Krause and Golovin, 2014).
+We leave this as future work.
+7 Limitations
+We provide an analysis of the impact of model size on the amenability to prompt optimization in Appendix A.1.
+However, in our evaluation, we only use GPT-3.5 (and in some cases GPT-4) as the solver LLM.
+We want to leave extensive evaluations of using open-source LLMs as solver LLMs, and perhaps even as expert LLMs, to future work.
+Further, we also want to evaluate on other generative tasks, besides the code generation task we study in the paper.
+Appendix A Appendix
+A.1 When is directional text optimization feasible?
+Figure 2: Estimating (probabilistic) Lipschitz constant of models (Definition 1) on (left) Ethos (middle) GSM8K and (right) MedQA datasets for GPT-4 and GPT-3.5 models.
+Consider the class of sequential algorithms like ProTeGi (Pryzant et al., 2023) and TextGrad (Hou et al., 2023) .
+The objective is to improve the accuracy of a given (black-box) solver LLM $f:\mathcal{X}\rightarrow\mathbb{R}$ that takes as input a prompt $\mathbf{x}\in\mathcal{X}$ and outputs the average accuracy on a validation set $D_{v}$.
+Since the set of prompts is combinatorially large, we assume that all prompts can be embedded in a vector space such that distance between two prompts in the space correspond to their semantic similarity.
+The prompt optimization problem can be written as $\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x};D_{v})$.
+Previous work has shown that LLMs can be brittle to their input: changing the prompt slightly can create a significant difference in performance (Zhuo et al., 2023).
+We want to understand if the optimization problem is well-conditioned.
+Typically, conditioning can be determined by the Hessian.
+However, since $f$ is black-box, we approximate it by measuring sensitivity, or more specifically, Lipschitz continuity near the optimal solution.
+Based on prior work on defining continuity of neural networks (Mangal et al., 2020), we use a probabilistic notion.
+Definition 1 (Probabilistic Lipschitz Continuity (Mangal et al., 2020)).
+Given a probability distribution over inputs $\mathcal{X}$, $r\geq 0$, and a distance measure $d$ such as $\ell_{1}$ or $\ell_{2}$ norm, a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(L,\epsilon)$-probabilistically Lipschitz with constant $L\geq 0$, if
+$\Pr_{\mathbf{x},\mathbf{x}^{\prime}\sim\mathcal{X}}[$ $\operatorname{d}(f(\mathbf{x}),f(\mathbf{x}^{\prime}))\leq L\cdot\operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})$
+$|\ \operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\leq r]\geq 1-\epsilon.$ (1)
+Note the focus on small changes in input through the parameter $r$.
+Intuitively, the Lipschitz property bounds the maximum change in $f$ given a small change in input prompt.
+Typically, the lower bound of error for any sequential optimization algorithm over $f$ is directly proportional to the Lipschitz constant $L$ (Malherbe and Vayatis, 2017).
+Therefore, for faster convergence, it is desirable to have a low $L$, especially near the optimal solution.
+Empirically, we estimate $L$ by sampling task-relevant prompts so that they are close to the optimal solution.
+Then we make small changes to the prompt such that the semantic meaning stays the same and measure the change in $f$ (See Appendix A.2 for experimental details).
+We show the change in $f$ per change in input for GPT4 and GPT3.5 models in Figure 2 for the Ethos, GSM8K and MedQA datasets.
+Assuming $\epsilon=0.05$, probabilistic Lipschitz constant $L$ for GPT4 is $<1$, whereas it is higher for GPT3.5.
+Thus, as the model sizes increases, the probabilistic Lipschitz constant decreases.
+So, larger models are more amenable to prompt optimization.
+A.2 Details on estimation of Lipschitz constant $L$
+TO calculate the Lipschitz constant for a given LLM and task, we take a human written promp and generate it’s paraphrases using GPT-4.
+We prompt GPT-4 with the following text: “You are given a sentence, you have to generate 30 paraphrases of the sentence, make sure that the core content of each paraphrase is same, you can use add, subtract or change words".
+These paraphrases are then evaluated on the validation set $D_{v}$.
+For a measure of distance between two prompts, we take the cosine similarity between the embeddings of two prompts.
+We use text-ada-002 for generating the text embeddings for prompts.
+A.3 Computational Complexity
+We now consider the compute complexity of UniPrompt in terms of the number of expert or solver LLM calls per epoch, stage-wise.
+Clustering: First, we evaluate all the training examples using the current prompt.
+Second, for every wrongly predicted example, we obtain feedback from the expert LLM.
+Third, for the given set of feedbacks, we use a single call to cluster it into $l$ clusters.
+Each of the above steps incurs $O(N)$ queries, so the total query complexity of the clustering stage is $O(N)$.
+Finally, for each example, i.e., (question, answer) pair, we simply map it to the $l$ clusters (no LLM calls).
+Mini-batch feedback and Batch-level aggregation: At a given epoch, we evaluate every question in the mini-batch using the current prompt and the solver LLM ($N$ queries overall).
+Next, we obtain one feedback over all the wrong questions in the mini-batch $m$ ($N/|m|$ queries).
+We use one call to aggregate these feedbacks.
+For prompt selection, we evaluate $4$ prompts on the batch $b$ (2 per beam), so $O(4|b|)$ queries per batch.
+Hence overall query complexity is $N+N/|m|+4N+1$ or $O(N)$.
+With LLM throughput of 0.5 qps, a training + validation set of 300 examples, 10 clusters, and 20 epochs, it takes under 7 hours to train.
+A.4 SLM Training Details
+To induce the ability of structured prompt generation in a smaller language model, we curate a section-wise dataset of around 12,000 task-prompt pairs.
+The tasks for training dataset creation were taken from tasksource library (Sileo, 2023) that contains around five hundred classification tasks.
+We extract the task description from tasksource-instruct, which contains tasksource dataset recasted with instructions.
+For instance, the task description for BIG-bench Entailed Polarity task is, "Given a fact, answer the following question with a yes or a no".
+The dataset provides diverse tasks and their short description, but not the human-generated prompts for each task.
+To approximate human-generated prompts, we use GPT-4 as a teacher model.
+By prompting GPT-4 with the task description and section description, we ask it to generate the contents of the section.
+To ensure that the generated section-wise prompts are concise and relevant, we prompt GPT-4 to not generate more than five lines of content for each section.
+We use LLAMA2-13B model, which we finetune using LoRA adapters as the auxiliary LM that generates sections.
+A.5 Data Set Creation of Real-World Task
 We sample real user queries from a proprietary application, rewrite them using ML models, and ask expert judges to label the query-pairs as identical or otherwise based on prescribed guidelines.
@@
 For testing, we use a separate labelled set of 2527 examples from two geographies — one where the training data was sampled from, and the other unseen.
-The prompt obtained using UniPrompt improves over the best manual prompt by 5.77% on the negative (rare) class, by 0.23% on the positive class, and by 1.86% overall on the test set.
-The learnt prompt captures the following facets of the task: (1) recognizing variations in names and abbreviations, and how they do not necessarily change the context; (2) recognizing the specificity of brands, and how even minor variations do change the context; and (3) recognizing the specificity of terms in queries, and how lack of specific terms can indicate departure of intent.
-4.4 Do diverse task facets organized as sections really help?
-We now revisit our hypothesis and motivation in Section 2.2.
-We want to empirically validate if all the diverse task facets that UniPrompt learns indeed contribute to the performance gains that we observe in Table 1.
-We consider two ablations:
-1) We successively remove each facet (i.e., sections) in the learnt prompt for the task and report the performances of the prompts with fewer facets.
-In Figure 3, for the Ethos dataset, we see that almost every additional facet contributes to non-trivial gains in accuracy.
-2) Could we have captured the information differently and retained the performance?
-We do a simple experiment – we summarize all the facets (i.e., learnt prompt) and evaluate the resulting prompt.
-In Figure 3 (right) (green line), we see that the summarized prompt has a significant accuracy drop.
-4.5 Ablations: Impact of Clustering, Inclusion of History, and Greedy Update
-Here we study effect of our design choices, namely clustering (Section 3.2), inclusion of history of past prompts (Section 3.2), and the greedy update rule (Section 3.2.3).
-The results are shown in Table 3.
-We see that clustering as well as edit history components are critical for performance of UniPrompt in all the datasets.
-We see a major drop of $14.8\%$ in accuracy in the Ethos dataset when clustering is removed, and a $4.3\%$ drop when history component is removed.
-In all the datasets except GSM8K, we find clustering is more important than history.
-This can attributed to variability of question types being less in GSM8K than in others: all questions are grade-8 arithmetic questions in GSM8K.
-We also find that the greedy update rule proves to be superior or competitive compared to beam search in relatively easier datasets — where even less exploration produces good results, greedy proves to be a more effective update rule.
-On the other hand, in more complex datasets like MedQA, greedy appears to be a bad strategy.
-We also see that clustering examples based on feedback (“Fb Clustering”) is a better strategy than clustering based on topics, except for the Ethos dataset.
-5 Related Work
-Here, we highlight relevant work that are not addressed in the manuscript so far.
-Deng et al.
-[6] present a discrete prompt optimization method, RLPrompt, using reinforcement learning, where a policy network learns to generate effective prompts through reward-based training, with an emphasis on enhancing training efficiency through effective reward stabilization techniques.
-A drawback of such automatic prompt optimization approaches [22, 36, 6, 31] is that the prompts generated tend to be short, often comprising only one or two sentences, which may not fully encapsulate the complexity of the task at hand.
-Prior research [28, 27] has highlighted the significance of specific sections within prompts.
-For instance, a well-designed prompt typically incorporates elements such as a system prompt and a chain-of-thought section.
-However, existing methods do not specifically target the optimization of individual sections and their respective contents within the prompts.
-Hsieh et al.
-[11] investigate the use of greedy and genetic algorithms to edit lengthy prompts.
-Their method focuses on paraphrasing one line at a time starting from an existing prompt, compared to our goal of learning facets of a task from scratch, thereby compromising generalization accuracy.
-Another orthogonal line of work that ha been explored recently for prompt optimization is the algorithmic selection of in-context examples [19, 9, 30, 25, 26].
-6 Limitations and Future Work
-We presented a method inspired by the human prompt engineering process to generate complex prompts from scratch that include different facets of a task.
-Our algorithm provides significant improvements over baseline prompt generation methods on multiple standard datasets.
-However, in our evaluation, we considered only classification tasks and only GPT-4 as the expert LLM.
-Generalizing to generative tasks, as well as investigating the impact of the expert LLM choice are some future directions.
-Appendix A Appendix
-Table 2: Comparison of UniPrompt (“Ours”) with MedPrompt, with GPT-4 as the solver model.
-Ours Ours + KNN Ours + KNN + COT Ours + KNN+COT+Ensemble MedPrompt
-MedQA 80.9 81.0 83.9 87.0 80.6
-PubMedQA 70.3 72.2 74.7 75.6 71.2
-MedMCQA 79.2 81.4 82.6 84.5 79.1
-MMLU MG 78.0 94.0 96.0 99.0 98.0
-Figure 3: Training curves for MedQA (left) and ARC (middle) datasets when UniPrompt is initialized with (published) state-of-the-art prompts; (right) ablation of facets on Ethos (Section 4.4).
-Table 3: Ablation of design choices in UniPrompt with GPT-3.5-Turbo as the solver model.
-Ethos ARC MedQA GSM8K
-UniPrompt -- History 88.0 84.6 55.3 80.8
-UniPrompt -- Clustering 77.5 82.0 54.1 81.5
-UniPrompt 92.3 86.0 57.1 82.4
-UniPrompt $+$ Greedy 93.7 90.5 55.5 82.3
-UniPrompt $+$ Fb Clustering 87.2 91.2 58.3 82.5
-A.1 Details on estimation of Lipschitz constant $L$
-TO calculate the Lipschitz constant for a given LLM and task, we take a human written promp and generate it’s paraphrases using GPT-4.
-We prompt GPT-4 with the following text: “You are given a sentence, you have to generate 30 paraphrases of the sentence, make sure that the cor content of each paraphrase is same, you can use add, subtract or change words".
-These paraphrases are then evaluated on the validation set $D_{v}$.
-For a measure of distance between two prompts, we take the cosine similarity between the embeddings of two prompts.
-We use text-ada-002 for generating the text embeddings for prompts.
-A.2 Details on marginal utility difference calculation for prompts
-For a given LLM and task, to study the marginal utility of different facets for a prompt, we first sample a few facets (generated using the fine-tuned model) that individually perform better than the chain-of-thought prompt as a baseline.
-Given these set of facets, we sample three facets $f_{1},f_{2},f_{3}$ and generate four prompts: $A:f_{1},B:f_{2},A+\Delta:f_{1}+f_{3},B+\Delta:f_{2}+f_{3}$.
-We define $\Delta_{A}$ as the accuracy difference between prompt $A$ and prompt $A+\Delta:f_{A}+f_{A+\Delta}$.
-Similarly, $\Delta_{B}=f_{B}+f_{B+\Delta}$
-A.3 Prompt Initialization
+A.6 Prompt to Llama2-13B for fine-tuning
+### Instruction:
+You are a prompt engineer, you have
+to write a structured prompt.
+For the given task description,
+examples and section description,
+write the contents of the section
+that align with
+section description.
+### Task Description:
+{data_point[’task_description’]}
+### Section Description:
+{data_point[’section’]}:
+{section_descriptions\
+[data_point[’section’]]}
+### Response:
+{data_point[’prompt’]}
+A.7 Prompt Initialization
 One line task descriptions:
@@
     MedQA: In this task, you are given a medical question. You have to solve the question.
+A.8 Additional Results
+Table 4: Performance (solved problems) of UniPrompt (GPT-4-Turbo solver) on code generation datasets, compared to GPT-4 (OpenAI et al., 2023) and newer models.
+Method HumanEval MBPP
+GPT-4 67.0 87.5
+GPT-4-Turbo 87.1 90.9
+GPT-4o 90.2 92.4
+UniPrompt 93.8 92.5
+Table 5: Ablation of LLM choices for UniPrompt on the Ethos dataset. ‘Init’ and ‘Final’ denote initial (i.e., task description) and final prompt accuracies.
+Expert LLM Solver LLM Init Final
+GPT-3.5-T GPT-3.5-T 76.8 82.4
+GPT-4 GPT-3.5-T 76.8 92.3
+GPT-3.5-T GPT-4 89.8 91.4
+GPT-4 GPT-4 89.8 94.3
+Table 6: Ablation of design choices in UniPrompt with GPT-3.5-Turbo as the solver model.
+Ethos ARC MedQA GSM8K
+UniPrompt -- History 88.0 84.6 55.3 80.8
+UniPrompt -- Clustering 77.5 82.0 54.1 81.5
+UniPrompt 92.3 86.0 57.1 82.4
+UniPrompt $+$ Greedy 93.7 90.5 55.5 82.3
+UniPrompt $+$ Fb Clustering 87.2 91.2 58.3 82.5
+Table 7: Ablation on the initial prompt for UniPrompt (best test accuracy in bold).
+Init Prompt Ethos ARC MedQA GSM8K
+Expert Prompt 84.0 86.0 52.3 82.4
+Llama Prompt 92.0 90.5 55.5 81.5
+Task Description 92.3 86.0 57.1 82.4
+Table 8: Comparison of UniPrompt (“Ours”) with MedPrompt, with GPT-4 as the solver model.
+MedQA PubMedQA MedMCQA MMLU MG
+Ours 80.9 70.3 79.2 78.0
+Ours + kNN 81.0 72.2 81.4 94.0
+Ours + kNN + CoT 83.9 74.7 82.6 96.0
+Ours + kNN+ CoT + Ensemble 87.0 75.6 84.5 99.0
+MedPrompt 80.6 71.2 79.1 98.0
 An example of sectioned initialization prompt generated using finetuned Llama Model
 Introduction:
-Assume the role of a science expert and answer the given
-question by selecting one of the options A, B, C or D.
-         1. Understand and solve science questions by selecting
-         the best answer from a given list of options.
-         2. Identify the logic behind the choices provided and
-         make an informed decision.
-         3. Use contextual clues to choose the most accurate
-         answer.
-         4. Be aware of the differences between science and
-         everyday language.
+Assume the role of a science expert and answer the given question by selecting one
+of the options A, B, C or D.
+     1. Understand and solve science questions by selecting the best answer
+     from a given list of options.
+     2. Identify the logic behind the choices provided and make an informed decision.
+     3. Use contextual clues to choose the most accurate answer.
+     4. Be aware of the differences between science and everyday language.
 Task Description:
@@
 answer options
-A.4 SLM Training Details
-To induce the ability of structured prompt generation in a smaller language model, we curate a section-wise dataset of around 12,000 task-prompt pairs.
-The tasks for training dataset creation were taken from tasksource library [24] that contains around five hundred classification tasks.
-We extract the task description from tasksource-instruct, which contains tasksource dataset recasted with instructions.
-For instance, the task description for BIG-bench Entailed Polarity task is, "Given a fact, answer the following question with a yes or a no".
-The dataset provides diverse tasks and their short description, but not the human-generated prompts for each task.
-To approximate human-generated prompts, we use GPT-4 as a teacher model.
-By prompting GPT-4 with the task description and section description, we ask it to generate the contents of the section.
-To ensure that the generated section-wise prompts are concise and relevant, we prompt GPT-4 to not generate more than five lines of content for each section.
-We use LLAMA2-13B model, which we finetune using LoRA adapters as the auxiliary LM that generates sections.
-A.5 Prompt to Llama2-13B for fine-tuning
-### Instruction:
-You are a prompt engineer, you have to write a structured prompt.
-For the given task description, examples and section description,
-write the contents of the section that align with section description.
-### Task Description:
-{data_point[’task_description’]}
-### Section Description:
-{data_point[’section’]}:{section_descriptions[data_point[’section’]]}
-### Response:
-{data_point[’prompt’]}
-A.5.1 Examples of training dataset
+A.8.1 Examples of training dataset
 Task: super glue boolq
@@
 be at the beginning or end of the sentence.
-A.6 Prompt for identifying important facets
+A.9 Prompt for identifying important facets
 you are given a task, along with it’s description, some examples
@@
 {sections}
-A.7 Clustering Type 1
+A.10 Clustering Type 1
 You are given a science question, you need to tell which broad
@@
 Subtopic:
-A.8 Clustering Type 2
+A.11 Clustering Type 2
 You are given a set of feedbacks, you need to cluster them into
@@
 between <Answer></Answer> tags like: <Answer>1</Answer>.’’’
-A.9 Feedback Prompts
+A.12 Feedback Prompts
 Feedback over mini-batch
-You are a teacher and you have to give feedback to your students on their answers.
+You are a teacher and you have to give feedback to your
+students on their answers.
 You are teaching how to solve math problems to your students.
-You are given a question, it’s true answer and answer given by student.
-You are also given the explanations written by your students while solving the questions.
+You are given a question, it’s true answer and answer given by
+student. You are also given the explanations written by your
+students while solving the questions.
 The questions are answered wrong by the students.
@@
 You are also given a set of wrong questions. You need to tell
-which edit can be applied to aid the student in solving the wrong question.
+which edit can be applied to aid the student in solving the
+wrong question.
 To achieve your task, try to follow the following steps;
@@
 delete_subsection(section_name, subsection_name)
-Your summary cannot include more than four functions. Make sure that the content is useful,
+Your summary cannot include more than four functions.
+Make sure that the content is useful,
 not just a very general statement. Something specific.
@@
 Summary:
-A.10 Editing Prompt
+A.13 Editing Prompt
 You are given an input prompt and a feedback, you have to
@@
 ### FINAL PROMPT
-A.11 Example of prompt evolution using our method
+A.14 Example of prompt evolution using our method
 See example in Figure 3.
@@
 After further iterations, our algorithm converges to a detailed, human-like longform prompt that achieves accuracy of $92\%$.
-A.12 Comparision of our method with existing methods
+A.15 Comparision of our method with existing methods
 See Figure 4.
@@
 This understanding is essential when dealing with problems that ask for the time since a past event or until a future event.
-Figure 6: Comparison of prompt produced by the state-of-the-art ORPO [31] and by UniPrompt on the GSM8K dataset.
-A.13 Effect of length on performance of prompt
+Figure 5: Comparison of prompt produced by the state-of-the-art ORPO (Yang et al., 2023) and by UniPrompt on the GSM8K dataset.
+A.16 Effect of length on performance of prompt
 Here we answer the question: How much does only length contribute to UniPrompt’s success?.
@@
 Summarized Prompt 84.3 85.5 66.0
+A.17 Do diverse task facets organized as sections really help?
+We want to empirically validate if all the diverse task facets that UniPrompt learns indeed contribute to the performance gains that we observe in Table 1. We consider two ablations:
+1) We successively remove each facet (i.e., sections) in the learnt prompt for the task and report the performances of the prompts with fewer facets.
+In Figure 6, for the Ethos dataset, we see that almost every additional facet contributes to non-trivial gains in accuracy.
+2) Could we have captured the information differently and retained the performance?
+We do a simple experiment – we summarize all the facets (i.e., learnt prompt) and evaluate the resulting prompt.
+In Figure 6 (right) (green line), we see that the summarized prompt has a significant accuracy drop.
+Table 10: Sensitivity of UniPrompt to expert LLM prompts, on the Ethos dataset.
+Expert LLM Prompt for UniPrompt Test Accuracy
+Simple prompt for mini-batch feedback 83.5
+Simple prompt for batch feedback 91.0
+Detailed prompts (Appendix A.12) 93.7
+A.18 Sensitivity to prompts used for expert LLMs in UniPrompt
+The prompts used for expert LLMs in our algorithm, i.e., for clustering, feedback over batches and mini-batches, and editing, do matter for obtaining good performance.
+However, note that the prompts are task-agnostic and can be used as-is for new tasks.
+Moreover, prompts for clustering and editing are very simple and involved minimal human effort.
+Further, to study the reliance of UniPrompt on the quality of feedback prompts, we run an ablation study, where we replace the engineered prompts for feedback at batch and mini-batch levels with simpler prompts.
+The results are given in Table 10 for the Ethos dataset.
+We observe that the performance of UniPrompt depends heavily on the prompt used for obtaining feedback at mini-batch level; whereas simplifying prompt for feedback at the batch level has much less impact on the final accuracy.
+A.19 UniPrompt training behavior
+An example of evolution of prompts using our algorithm is given in Appendix 3.
+It starts with a simple description of task and adds important facets like differentiating between hate speech and rudeness.
+In contrast, ProTeGi (Pryzant et al., 2023) yields a rather terse prompt on the same dataset: “Does the following text contain language that targets a group of people based on their religion, gender, or other personal characteristics?”.
+Figure 6: Training curves for MedQA (left) and ARC (middle) datasets when UniPrompt is initialized with (published) state-of-the-art prompts; (right) ablation of facets on Ethos.
+The training curves in Figure 6 show that our method initially performs edits on the prompt that simultaneously increase the train as well as the validation accuracy.
+After about 10 or 15 iterations (each batch update is an iteration), validation accuracy decreases while train accuracy continues increasing, indicating overfitting; which we overcome using early stopping.